### v1

# ThinK: Thinner Key Cache by Query-Driven Pruning

Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications by leveraging increased model sizes and sequence lengths. However, the associated rise in computational and memory costs poses significant challenges, particularly in managing long sequences due to the quadratic complexity of the transformer attention mechanism. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence lengths, we uncover that the channel dimension of the KV cache exhibits significant redundancy, characterized by unbalanced magnitude distribution and low-rank structure in attention weights. Based on these observations, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in memory costs by over 20% compared with vanilla KV cache eviction methods. Extensive evaluations on the LLaMA3 and Mistral models across various long-sequence datasets confirm the efficacy of ThinK, setting a new precedent for efficient LLM deployment without compromising performance. We also outline the potential of extending our method to value cache pruning, demonstrating ThinK’s versatility and broad applicability in reducing both memory and computational overheads.
1 Introduction
Large language models (LLMs) (Hadi et al.,, 2023; Brown et al.,, 2020; OpenAI,, 2023; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Scao et al.,, 2022; Reid et al.,, 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks. A key principle, the Scaling Law (Kaplan et al.,, 2020), suggests that LLMs exhibit emergent abilities as model size increases, enhancing their capacity to understand context and handle long sequences (Xiong et al.,, 2023). This capacity growth allows LLMs to generate coherent and contextually accurate responses and enables various downstream applications, such as document summarization (Zhang et al.,, 2019; Zhang et al., 2024a, ), code generation (Chen et al., 2021b, ), and conversational AI (Bordes et al.,, 2016; OpenAI,, 2022), .
Despite their success in various applications, the generation of LLMs incurs significant expenses, which escalate with increasing model size and sequence length. Notably, both the training (Strubell et al.,, 2020; Hoffmann et al.,, 2022; Dong et al., 2024a, ) and inference (Ainslie et al.,, 2023) stages involve frequent generation by LLMs, further contributing to these costs. Consequently, efficient LLMs have gained popularity in recent years (Hu et al.,, 2021; Wan et al.,, 2023). To address these challenges, quantization (Frantar et al.,, 2022; Lin et al.,, 2024; Dettmers et al.,, 2024; Xu et al.,, 2023) and pruning methods (Frankle and Carbin,, 2018; Blalock et al.,, 2020) are employed to reduce model size. Additionally, managing long sequences presents another cost due to the transformer attention mechanism. The quadratic complexity of the attention mechanism results in substantial computational burden when dealing with long sequences, scaling poorly with sequence length. Therefore, effective management of long sequences is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario and aim to reduce the memory consumption associated with lengthy sequences.
Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size of each head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference. To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or performing quantization over the caches. It is well-acknowledged that token importance tends to be sparse. Consequently, KV eviction algorithms have been proposed to minimize the KV cache memory cost from dimension $S$ (Xiao et al., 2023b, ; Li et al.,, 2024; Zhang et al., 2024c, ; Leviathan et al.,, 2023). Additionally, inter-layer redundancy has been explored (Liu et al., 2024a, ; Wu and Tu,, 2024; Brandon et al.,, 2024) to address the layer dimension $L$. Despite these advances, existing methods have largely overlooked the channel dimension $D$. In this paper, we highlight that the magnitude across Key cache channel dimensions is significantly unbalanced. Additionally, we observe that the attention weights exhibit a low-rank structure. Based on these findings, we hypothesize that the channel dimension of the key cache exhibits redundancy. Consequently, we explore the redundancy of KV cache tensors specifically from dimension $D$, aiming to devise effective strategies that reduce costs without compromising performance.
In our paper, we propose a simple yet effective method – ThinK, for KV cache pruning. To pinpoint the least significant channels, we formulate the task as an optimization problem, aiming to minimize the loss in attention weights attributable to pruning. To effectively address this problem, we establish a novel query-dependent criterion that assesses the importance of each channel. Using this criterion, we then select the most critical channels in a greedy fashion. We evaluate ThinK using the LLaMA3 (Meta,, 2024) and Mistral (Jiang et al.,, 2023) models and validate its effectiveness across various long-sequence datasets. The results indicate that when paired with token eviction methods, ThinK not only achieves comparable or superior accuracy but also reduces KV cache memory costs by more than 20%.
Contributions.
This work pioneers the investigation into the sparsity structure of the channels in the KV cache. Specifically, we discovered that the activated key cache is sparse given a specific query. This insight allows us to prune the key cache channels using a query-induced norm. Inspired by this phenomenon, we introduce a plug-and-play method ThinK, the first channel pruning method specifically designed for KV cache. ThinK reduces the dimensionality of the cache channel, leading to linear savings in both memory and computational requirements. As plug-and-play techniques, our proposed ThinK is orthogonal to other KV cache compression schemes (e.g. KV cache eviction, quantization). Our extensive experiments demonstrate the remarkable efficiency of ThinK on LLaMA-3 and Mistral models. Furthermore, we explore the promising potential of extending ThinK to value cache pruning (ThinKV) in specific scenarios, highlighting the broad applicability of our method.
Figure 1: An illustration of the pruning procedure of ThinK. Within each head, scores are calculated for each channel, and only the top T channels out of D are selected for retention. A binary channel mask, along with the pruned keys, is subsequently stored in the cache memory.
2 Observations
We identify several key observations that motivate our approach to prune the channel of KV-Cache. Specifically, we visualize the magnitude of KV-cache and perform singular value decomposition of the attention in LLaMA3-8B model. We show the visualization for certain heads and layers for illustration purposes.
Magnitude of KV-Cache Channel
Figure 2 visualizes the absolute values of key and value cache across the tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis.. Consistent with previous findings (Lin et al.,, 2024; Xiao et al., 2023a, ; Liu et al., 2024b, ), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns. For instance, the absolute value around $50^{th}$ channel of key cache has much larger magnitudes than other channels for all tokens in layer $14$ (Figure 2 (a)). The same observation applies to $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 2 (c)). Given such an observation, Liu et al., 2024b proposed to perform quantization over the channels of the key cache based on such observations. In addition to channel quantization, the observation also suggests that it is feasible to prune away certain key cache channels that contribute less to the attention mechanism. Moreover, channel quantization and pruning techniques are orthogonal such that they can be utilized concurrently to enhance model efficiency.
(a) Key Cache ($14^{th}$ layer)
(b) Value Cache ($14^{th}$ layer)
(c) Key Cache ($20^{th}$ layer)
(d) Value Cache ($20^{th}$ layer)
Figure 2: Magnitude of key and value cache for LLaMA3-8B. The first head of layer 14 and layer 20 of LLaMA3-8B is selected to visualize the magnitude of the key and value caches. We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
Singular Value Analysis
We conducted singular value decomposition (SVD) (Demmel,, 1997) on the attention of the specified layers to investigate their principal components. The singular values derived from SVD capture the effective rank of the attention matrix, indicating how the information is distributed across different components.
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}$ $=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}$ (1)
Figure 3 (a) illustrates the energy distribution of the singular values, plotted in log scale to enhance the visibility of differences among them. Notably, only a few singular values exhibit high energy levels exceeding $0.01$ across all heads and layers, highlighting their relative significance. This observation aligns with previous findings (Bhojanapalli et al.,, 2021), where a small subset of singular values often captures most of the information in attention mechanisms. In addition, the rapid decay of the energy suggests that a low-rank approximation can effectively capture the essential information in the key cache.
Figure 3 (b), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy. These findings suggest that the attention matrix is inherently low-rank (Wang et al.,, 2020; Chen et al., 2021a, ; Dong et al., 2024b, ), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al.,, 2024).
(a) Energy of singular values
(b) Cumulative energy of singular values
Figure 3: The energy and cumulative energy of the singular values.
3 ThinK
Notations.
[…]
3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows:
$2\times B\times S\times L\times N\times D,$ (2)
where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cach size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al.,, 2017), such as LLaMA2-7B (Touvron et al., 2023b, ), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al.,, 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b, ; Zhang et al., 2024c, ; Li et al.,, 2024) and precision (Hooper et al.,, 2024; Liu et al., 2024b, ). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
Table 1: Performance comparison of pruning K cache by $l_{p}$ norm on LongBench.
Method $\lambda$ Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O 0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+$l_{1}$ 30 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
+$l_{1}$ 40 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
+$l_{2}$ 30 23.98 17.04 35.19 39.27 31.29 20.4 21.62 22.46 23.34 40.5 89.75 40.71 5.54 68.67 60.12 58.52 37.40
+$l_{2}$ 40 23.76 16.23 32.19 40.23 32.13 20.69 21.3 22.25 23.2 39.5 89.61 40.24 5.66 69.0 60.09 59.45 37.22
SnapKV 0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
+$l_{1}$ 30 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
+$l_{1}$ 40 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
+$l_{2}$ 30 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.5 90.33 40.31 5.7 68.42 62.65 60.07 40.28
+$l_{2}$ 40 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
Magnitude base Pruning: Based on the observations in Figure 2 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
$M_{n,d}=\big{\|}K[n,:,d]\big{\|}_{p}.$ (3)
Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head:
$I=\textbf{Top}_{T}(M,T)$ (4)
where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel. $n\in[1,N]$ and $d\in[1,D]$ are indicators of heads and channels in key cache. $I\in\mathbb{(Z^{+})}^{N\times T}$ stores the indicators of the top $T$ values in tensor $M$ per head.
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al.,, 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c, ) and SnapKV (Li et al.,, 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.
3.2 Query-Driven Pruning
For each head, the attention scores are computed using the queries and keys, and then applied to the values. The formula for the attention for head $i$ is:
$\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}$ (5)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. When pruning one channels of $\textbf{K}_{i}$, the corresponding channels in $\textbf{Q}_{i}$ will also be removed. We aim to find the optimal subset of channels to prune, denoted by the selection matrix $\textbf{S}\in\{0,1\}^{D\times D}$, where S is a diagonal matrix with binary entries ($1$ for keeping a channel, 00 for pruning it). To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights:
$\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}$ (6)
Given a pruning ratio $\lambda$, it can further expanded as:
$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$
subject to $\quad\text{trace}(\mathbf{S})=\lfloor(1-\lambda)D\rfloor$
$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$ (7)
For simplicity, we use greedy algorithm to optimize $\mathbf{S}$. To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores:
$\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F}$ (8)
$I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$ (9)
Here’s a detailed explanation of why it optimizes the selection matrix. The $\text{score}_{i}[j]$ measures the magnitude of the interaction between the query and key vectors for channel $j$ in each head $i$. By selecting channels with the highest interaction magnitudes, we aim to retain the most significant contributions to the attention mechanism. This criterion ensures that the selected channels preserve the primary information flow in the attention computation, thereby minimizing the loss of important information.
Observation Window. Based on the observations in SnapKV (Li et al.,, 2024) that the last window of input sequence recognizes highly similar attention allocation pattern with the actual generation. To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
Figure 4: Implementation during decoding.
3.3 Implementations
Following SnapKV, we focus on the long context input scenario. We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size. Additionally, we maintain a binary mask to indicate which channels have been pruned. Note that the memory overhead associated with this mask is negligible. Figure 4 illustrates one implementation of our method during the decoding stage. The pruned keys are first zero-filled using the mask to restore them to their original size, and then concatenated with the unpruned keys. This implementation can be effectively integrated with optimization techniques such as FlashAttention (Dao,, 2023). The second implementation involves initially pruning the Query using the mask. The pruned Query is then multiplied by the pruned Key, while the unpruned Query is applied to the unpruned Key. Subsequently, the two outputs are concatenated. In theory, this implementation has lower computational costs.
4 Experiment Results
In this section, we conduct extensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1. Next, we report the performance of our KV pruning method with two different KV compression methods in long-context scenarios in Section 4.2. Finally, we present the results of ablation experiments in Section 4.4.
4.1 Settings
Benchmark Datasets.
We perform evaluations of our method with state-of-the-art KV cache compression methods on two widely used benchmarks: LongBench and Needle-in-a-Haystack. LongBench (Bai et al.,, 2023) is designed to comprehensively evaluate LLM’s long context understanding capabilities. It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference. Needle-in-a-Haystack (Kamradt,, 2023) is a recent popular test challenge that requires models to accurately identify a small piece of information (“needle”) in a long document (“haystack”), where the needle is placed at a random position. This challenge can test if KV cache compression methods still retain the small piece of critical information.
Baseline Approaches.
The baseline methods include Heavy Hitter Oracle (H2O) and SnapKV, both of which are the state-of-the-art KV cache compression methods but use different strategies for selecting KV cache. H2O (Zhang et al., 2024c, ) is designed to reduce memory usage by dynamically balancing recent tokens and Heavy Hitter (H2) tokens, where H2 tokens are a small set of tokens that contribute most of the value when computing attention scores. SnapKV (Li et al.,, 2024) automatically compresses KV caches by selecting clustered important KV positions for each attention head. This is inspired by the fact that each attention head has a specific pattern, focusing on certain attention features during generation.
Table 2: Performance comparison of Key cache pruning on LLaMA3-8B by ThinK on LongBench.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
KV-size Full
ALL KV 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
KV-size 128
H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
+ThinK(40) 22.85 14.55 29.49 38.63 30.84 18.90 20.12 21.96 20.68 38.50 86.38 38.40 5.50 69.17 57.93 56.12 35.63
+ThinK(50) 23.47 14.06 28.67 38.35 30.21 17.87 19.69 21.94 19.95 38.50 87.14 38.07 4.92 69.50 57.99 56.66 35.44
[…]
+ThinK(40) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
+ThinK(50) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
KV-size 512
H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+ThinK(40) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
+ThinK(50) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
+ThinK(60) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
[…]
+ThinK(40) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
+ThinK(50) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
+ThinK(60) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
KV-size 1024
H2O 25.62 22.16 36.81 41.01 33.53 19.41 23.28 22.65 25.41 46.50 90.82 41.78 5.79 69.25 59.69 55.50 38.70
+ThinK(40) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
+ThinK(50) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
+ThinK(60) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
[…]
+ThinK(40) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
+ThinK(50) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
+ThinK(60) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
KV-size 2048
H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
+ThinK(40) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
+ThinK(50) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
+ThinK(60) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
[…]
+ThinK(40) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
+ThinK(50) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
+ThinK(60) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
Implementation Details.
In this paper, we use LLaMA3-8B-Instruct (Meta,, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al.,, 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al.,, 2020). Our ThinK aims to prune channels of the KV, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using one Nvidia A100. To fairly compare KV cache compression methods and KV cache compression integrated with ThinK, we use the same hyperparameters for both. For example, when comparing SnapKV and SnapKV integrated with ThinK, we set the maximum pooling kernel size to 7 and the observation window size to 32, using the same KV-size for both.
Table 3: Performance comparison of Key cache pruning on Mistral-7B by ThinK on LongBench.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
KV-size Full
ALL KV 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
KV-size 128
H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
+ThinK(40) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
+ThinK(50) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
+ThinK(60) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
[…]
+ThinK(40) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
+ThinK(50) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
+ThinK(60) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
KV-size 512
H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
+ThinK(40) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
+ThinK(50) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
+ThinK(60) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
[…]
+ThinK(40) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
+ThinK(50) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
+ThinK(60) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
KV-size 1024
H2O 23.67 28.55 46.40 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
+ThinK(40) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
+ThinK(50) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
+ThinK(60) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
[…]
+ThinK(40) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
+ThinK(50) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
+ThinK(60) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
KV-size 2048
H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.00 86.35 42.48 2.72 86.64 56.98 53.91 40.69
+ThinK(40) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
+ThinK(50) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
+ThinK(60) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
[…]
+ThinK(40) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
+ThinK(50) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
+ThinK(60) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52
4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and KV compression methods integrated with our channel pruning method of Key cache (ThinK) over two different base LLMs across various KV-sizes on LongBench. The following observations can be drawn: (1) Our method can further prune the channels of the key cache after compressing the KV cache with H2O and SnapKV. For the base model LLaMA3-8B, our method reduces memory usage and slightly improves performance for both H2O and SnapKV. For the base model Mistral-7B, our method reduces memory with only a slight drop in performance in some cases. (2) Comparing the results of SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm shown in Table 1, our proposed query-based channel pruning shows better performance when the pruning ratio is $40\%$ compared to no pruning. However, $l_{1}$ and $l_{2}$ norm methods cannot outperform the KV compression methods without pruning. (3) When the pruning ratio is small, performance tends to be better than when the pruning ratio is large. (4) When the KV-size is increased from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV cache size of $2048$ and a pruning ratio of $40\%$, our method can even outperform the LLaMA3-8B with a full KV cache. The above observations indicate that our method is agnostic to existing KV cache compression methods and can further improve their performance and memory reduction. Additionally, query-based pruning is more effective than $l_{1}$ and $l_{2}$ norm for channel pruning in LLMs.
4.3 Results on Needle-in-a-Haystack
Table 4 presents the results of the Needle-in-a-Haystack test based on the SnapKV (Li et al.,, 2024) approach with KV cache size of input prompt from $128$ to $2048$. With a modest pruning ratio $\lambda=40$%, the accuracy of ThinK is consistently better or comparable than the original SnapKV across both LLaMA3 and Mistral with different KV cache sizes. Such comparisons demonstrate that the proposed query-driven channel pruning can effectively retain the informative channels and discard the noisy ones. With a pruning ratio of $\lambda\geq 50$%, we observe an accuracy drop with a small KV-cache size, particularly for $128$ and $512$ across both LLaMA3 and Mistral. However, ThinK is still able to obtain comparable performance with SnapKV when we have a larger KV cache size (i.e., $1024$ and $2048$). Intuitively, a larger pruning ratio with a smaller KV cache size could tend to lose more important information compared to a larger KV cache size. In addition, the performance on larger KV cache size also suggests ThinK is robust for long context retrieval tasks.
Figure 6 (a)-(d) visualize the Needle-in-a-Haystack test accuracy against different token lengths and depth percent. The KV cache size is configured to $128$ and $1024$, with pruning ratio $\lambda$ equal to $40$% and $50$%, respectively. First, ThinK can maintain the original retrieval capability of SnapKV though the overall numerical accuracy is different (i.e., $77.8$ vs $78.6$ and $90.4$ vs $90.6$). Because ThinKachieves the same accuracy at most of the entries with different token limits and depths as Snap KV. In addition, ThinK also successfully retrieves some “needles” where SnapKV fails, leading to higher overall accuracy. The visualization results further demonstrate the robustness of ThinK in a fine-grained perspective and ThinK can enhance the original quantization approach.
Table 4: Needles-in-a-Haystack Test Results
Model Method $\bm{\lambda}$ KV-size
128 512 1024 2048
LLaMA3-8B SnapKV 00 79.6 90.2 91.2 91.7
SnapKV+ThinK 40 79.6 90.3 91.2 91.7
SnapKV+ThinK 50 77.4 89.6 91.0 91.7
Mistral-7B SnapKV 00 77.8 89.5 90.4 90.8
SnapKV+ThinK 40 78.6 90.1 90.6 90.9
SnapKV+ThinK 50 78.1 90.1 90.8 91.1
SnapKV+ThinK 60 75.9 89.2 90.6 91.1
4.4 Ablation Studies
Table 5: Performance comparison of pruning K cache using different recent-size.
Recent-Size Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O + ThinK($\lambda=40\%$)
0 24.40 27.50 45.42 35.17 24.45 13.02 27.65 23.88 26.86 53.50 86.06 41.73 3.01 83.42 55.12 51.32 38.91
32 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
128 25.69 30.93 48.32 39.63 26.08 16.82 27.18 23.92 26.62 53.50 86.39 42.96 3.29 86.39 56.77 53.60 40.51
SnapKV + ThinK($\lambda=40\%$)
0 24.94 28.58 45.78 39.59 25.40 15.92 29.50 24.05 26.72 70.00 85.60 41.38 2.97 84.00 55.27 52.39 40.76
32 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
128 25.75 32.49 48.61 41.01 27.18 19.14 28.79 24.64 26.77 70.00 86.37 42.77 3.61 87.13 57.19 54.39 42.24
Table 6: ablations with the same memory consumption.
Methods Memory(M) Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
H2O
Vanilla 54.5 21.29 20.69 37.66 28.65 21.08 14.01 20.20 22.11 21.33 38.50 82.55 39.87 3.66 78.14 50.32 48.54 34.29
ThinK 54.4 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
Vanilla 208.0 22.13 23.83 43.24 30.92 23.36 14.56 22.92 22.77 24.23 41.50 85.04 41.26 3.02 86.03 54.91 50.50 36.89
ThinK 208.0 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
Vanilla 413.0 22.90 28.45 46.16 35.57 23.86 13.74 24.90 23.19 25.77 44.50 85.54 41.97 3.22 85.82 55.96 52.33 38.37
ThinK 412.8 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
Vanilla 822.5 25.51 30.23 48.23 39.72 25.56 16.75 26.98 23.81 26.47 50.50 86.43 42.09 2.78 85.57 57.4 53.42 40.09
ThinK 822.4 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
SnapKV
Vanilla 54.5 19.25 19.95 42.80 35.88 21.96 14.59 18.76 21.71 20.70 46.00 84.12 39.43 2.59 65.36 51.39 47.81 34.52
ThinK 54.4 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
Vanilla 208.0 23.31 27.45 48.85 38.77 23.93 16.50 23.44 23.63 24.13 66.00 86.05 41.00 2.62 87.01 56.13 52.60 40.09
ThinK 208.0 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.90 42.51 2.92 85.32 55.89 53.35 40.40
Vanilla 413.0 24.24 29.53 49.13 40.48 25.05 18.74 25.46 23.64 25.60 68.00 86.14 41.42 3.03 88.55 57.08 53.86 41.25
ThinK 412.8 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.30 56.71 54.19 41.62
Vanilla 822.5 24.84 31.90 48.16 41.32 26.77 19.49 28.23 24.63 26.41 70.00 86.32 41.83 2.91 88.06 56.98 53.74 41.97
ThinK 822.4 25.77 32.67 48.7 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
Impact of Different Recent Sizes.
Preserve the most recent KV embeddings (Zhang et al., 2024c, ; Li et al.,, 2024) is important for maintaining the performance of LLMs after KV cache compression. Note that a tradeoff exists: increasing the recent-size allows more infomation to be propagated, while increasing the cache size to be stored. To study its impacts, we evaluate the performance produced by three recent-size, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the basleline model. The results are summarized in Table 5. One can observe that a recent-size of 32 yields superior performance than 00 in terms of averaged score on LongBench which demonstrates the importance of keeping the most recent KVs. On the other hand, the performance of $32$ and $128$ is negligible, suggesting that maintaining the most recent 32 KVs suffices to preserve optimal performance.
Figure 5: Performance comparisons under the same KV cache memory consumption.
Comparison Under the Same Memory Consumption.
To ensure a fair comparison, we adjust the KV cache size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK. For example, the KV cache size of H2O with ThinK is $128$. Due to channel pruning of the key cache, the memory consumption of H2O with ThinK at a $128$ KV-size is lower than H2O at a 128 KV-size. Therefore, the KV-size of H2O is adjusted from $128$ to $109$. Table 6 and Figure 5 present the results of these comparisons on LongBench. As the results show, H2O or SnapKV with ThinK deliver superior performance compared to H2O or SnapKV without ThinK while maintaining the same memory usage. This demonstrates that KV cache compression methods, combined with our query-based channel pruning, can compress the KV cache more effectively and utilize memory more efficiently.
Pruning Channels of Both Key and Value Cache.
In our previous study, we evaluate methods with pruning channels of the key cache. In this part, we explore the impact of pruning channels of the value cache (Appendix A.1). Specifically, for KV cache compression methods, we prune channels of the key cache at one pruning ratio and channels of the value cache at another pruning ratio. Table 7 presents the results with LLaMA3-8B and Mistral-7B on LongBench. When testing on the base model LLaMA3-8B, H2O or SnapKV with key and value channel pruning can perform on par with H2O or SnapKV without channel pruning. In some cases, H2O or SnapKV with key and value channel pruning may even perform better. Pruning the channels of the value cache over the base model Mistral-7B experiences a slight performance drop. Despite this, note that KV cache with additional value channel pruning can further reduce memory usage.
Table 7: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench. H2O $+$ ThinK($\lambda_{1}$+$\lambda_{2}$) indicates that H2O has $\lambda_{1}$of the Key cache channels pruned and $\lambda_{2}$of the Value cache channels pruned.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
LLaMA3 8B KV-size 128
H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
+ThinK(30+30) 23.71 13.65 33.08 41.86 29.88 18.04 19.60 21.65 20.26 38.00 86.08 38.61 5.16 69.50 57.59 55.19 35.74
[…]
+ThinK(30+30) 21.86 13.79 33.26 40.93 29.39 19.22 18.81 21.30 19.26 41.50 87.00 37.95 5.78 69.50 57.84 55.62 35.81
KV-size 512
H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
+ThinK(30+30) 22.83 17.57 34.18 42.67 33.52 19.95 21.17 22.23 22.82 38.50 90.11 39.08 5.21 69.0 59.99 56.83 37.23
[…]
+ThinK(30+30) 24.57 24.59 38.09 44.61 34.37 20.37 21.23 21.95 23.30 66.00 90.69 39.38 5.60 69.00 61.75 58.46 40.25
KV-size 2048
H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
+ThinK(30+30) 25.03 26.77 39.68 42.12 33.08 19.59 23.00 22.89 25.27 51.00 91.11 40.58 5.23 69.00 61.12 57.95 39.59
+ThinK(40+40) 24.87 24.31 37.77 43.13 34.42 19.60 21.67 22.70 24.52 49.00 90.81 39.28 6.00 69.00 61.81 58.08 39.19
[…]
+ThinK(30+30) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
+ThinK(40+40) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
Mistral 7B KV-size 128
H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
+ThinK(30+30) 20.71 21.49 38.01 30.66 22.28 13.87 20.13 22.45 21.07 38.50 82.20 38.69 2.94 78.56 51.55 48.28 34.46
[…]
+ThinK(30+30) 19.92 20.61 42.68 37.63 23.19 15.09 18.97 21.93 20.55 45.00 84.06 39.33 2.99 66.00 51.51 47.51 34.81
KV-size 512
H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
+ThinK(30+30) 22.36 24.26 44.77 30.47 22.94 14.96 22.63 22.90 23.73 41.50 85.30 40.21 3.08 80.07 54.48 50.96 36.54
+ThinK(30+10) 22.14 25.15 45.29 31.78 23.21 14.62 23.36 22.70 24.51 41.50 85.61 41.58 2.75 84.03 54.50 51.09 37.11
[…]
+ThinK(30+30) 24.10 27.04 47.76 38.66 25.45 17.51 22.64 22.81 23.91 66.00 86.62 39.91 3.36 82.24 55.96 52.81 39.80
+ThinK(30+10) 23.90 28.14 48.35 39.03 24.83 16.68 23.51 23.12 24.34 67.50 86.09 41.69 2.65 84.34 57.29 53.22 40.29
KV-size 1024
H2O 23.67 28.55 46.4 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
+ThinK(30+30) 23.65 26.54 47.00 35.52 24.79 17.15 23.64 23.12 25.20 44.00 86.38 41.67 3.46 80.14 56.53 52.86 38.23
+ThinK(30+10) 24.13 28.57 46.31 35.59 24.92 15.34 24.58 23.33 25.93 45.50 85.91 42.97 2.57 83.64 55.39 52.73 38.59
[…]
+ThinK(30+30) 25.29 29.25 49.17 41.25 25.75 19.37 24.64 23.02 25.27 69.00 86.70 40.92 3.29 82.06 57.15 54.15 41.02
+ThinK(30+10) 25.84 29.30 49.56 41.44 25.29 19.02 25.21 23.73 25.72 69.00 86.69 42.55 2.44 85.76 57.55 54.10 41.45
KV-size 2048
H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.0 86.35 42.48 2.72 86.64 56.98 53.91 40.69
+ThinK(30+30) 25.60 28.74 47.54 38.67 26.25 17.35 24.54 23.27 26.15 51.00 87.01 43.02 2.94 81.46 56.41 54.26 39.64
+ThinK(30+10) 25.64 30.65 48.95 40.42 26.43 16.65 26.76 23.51 26.59 52.50 86.53 43.45 2.66 83.96 56.55 53.83 40.32
[…]
+ThinK(30+30) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
+ThinK(30+10) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94
(a) KV-size 128, SnapKV, Acc.77.8
(b) KV-size 128, SnapKV+ThinK($40$) Acc.78.6
(c) KV-size 1024, SnapKV, Acc.90.4
(d) KV-size 1024, SnapKV+ThinK($50$), Acc.90.8
Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2.
5 Related Work
In scenarios involving long contexts, the most significant computational burden from the attention mechanism is the key-value (KV) cache. Reducing the KV cache is a high priority for optimizing deployment efficiency. System-level optimizations, such as FlashAttention (Dao,, 2023) and PagedAttention (Kwon et al.,, 2023), have been developed to address this issue. Additionally, algorithm-level optimizations are being explored to further enhance efficiency.
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b, ) maintains a few initial tokens and some recent tokens based on the observation of attention sink, which may result in the loss of important information carried by the dropped tokens. H2O (Zhang et al., 2024c, ) retains only a small portion of the tokens by greedily dropping tokens based on their contributions to the cumulative attention. SnapKV (Li et al.,, 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al.,, 2023) adaptively evicts tokens from attention heads that emphasize local contexts. This approach focuses on discarding non-special tokens centered on special tokens, while the standard KV cache is used only for attention heads that broadly attend to all tokens. PyramidKV (Zhang et al., 2024b, ) and PyramidInfer (Yang et al.,, 2024) considers adjusting the KV cache size across different layers by allocating more cache in the lower layers and less in the higher ones.
KV Cache Quantization. SmoothQuant (Xiao et al., 2023a, ) can quantize the KV cache to 8-bit with minimal performance degradation. Q-Hitter (Zhang et al., 2024c, ) uses accumulated attention scores and ‘Quantization Friendliness’ metrics to identify tokens that are essential for maintaining the generalization capabilities of LLMs and are suitable for KV cache quantization. Some studies have found that the key cache and value cache should be quantized differently (Liu et al., 2024b, ; Hooper et al.,, 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
Structured Pruning of LLMs. Structured pruning  (Ma et al.,, 2023; Ding et al.,, 2023) of LLMs traditionally focuses on removing unimportant layers, heads, and hidden dimensions, which often results in significant performance degradation. In contrast, our methodology preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels based on data dependant criterion, our approach greatly reduce the key cache-size with negligible performance loss.
6 End Note and Future Direction
Motivated by the observation that certain channels have significantly larger magnitudes compared to others, and the singular value analysis indicates that the key cache is inherently low-rank (Singhania et al.,, 2024), we propose ThinK to perform pruning over the key cache channel. The proposed pruning strategy is query-dependant and optimized based on the attention scores, ensuring that essential information is retained for each input query. In addition, ThinK can be seamlessly integrated with other popular token-level KV cache quantization techniques (Li et al.,, 2024; Zhang et al., 2024c, ), further enhancing inference efficiency. Extensive experiments on LongBench (Bai et al.,, 2023) and Needle-in-a-Haystack tests with two foundation models demonstrate the effectiveness and robustness of our query-dependent channel pruning method. Our approach achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%. Our analysis indicates that ThinK can maintain superior performance over baselines with a smaller KV cache size under equivalent memory consumption conditions.
Future work will focus on strategies to increase the pruning ratio without compromising model performance, and exploring pruning techniques for value caches. We also plan to evaluate channel importance from a more sophisticated and compositional perspective, incorporating both token-level and channel-level information to calculate importance.
Appendix A Additional Experimental Details
A.1 Value Cache Pruning
Similar to the approach used for the Key cache, the pruning of channels in the Value cache can be guided by two primary criteria: magnitude-based pruning and query-driven pruning. We find that query-driven pruning is still better than magnitude based pruning.
$\text{Score}_{v,i}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})[j]=\|\text{softmax}(\frac{\textbf{Q}_{i}[-S_{\text{obs}}:]\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}[:,j]\|_{F}$ (10)
$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (11)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache. Then, only top T channels are retained.

### v2

@@
 Abstract
-Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications by leveraging increased model sizes and sequence lengths.
-However, the associated rise in computational and memory costs poses significant challenges, particularly in managing long sequences due to the quadratic complexity of the transformer attention mechanism.
+Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications.
+However, their increased computational and memory demands present significant challenges, especially when handling long sequences.
 This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference.
-Unlike existing approaches that optimize the memory based on the sequence lengths, we uncover that the channel dimension of the KV cache exhibits significant redundancy, characterized by unbalanced magnitude distribution and low-rank structure in attention weights.
-Based on these observations, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels.
-Our approach not only maintains or enhances model accuracy but also achieves a reduction in memory costs by over 20% compared with vanilla KV cache eviction methods.
-Extensive evaluations on the LLaMA3 and Mistral models across various long-sequence datasets confirm the efficacy of ThinK, setting a new precedent for efficient LLM deployment without compromising performance.
-We also outline the potential of extending our method to value cache pruning, demonstrating ThinK’s versatility and broad applicability in reducing both memory and computational overheads.
+Unlike existing approaches that optimize the memory based on the sequence length, we identify substantial redundancy in the channel dimension of the KV cache, as indicated by an uneven magnitude distribution and a low-rank structure in the attention weights.
+In response, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels.
+Our approach not only maintains or enhances model accuracy but also achieves a reduction in KV cache memory costs by over 20% compared with vanilla KV cache eviction and quantization methods.
+For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.
+Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK, establishing a new baseline algorithm for efficient LLM deployment without compromising performance.
 1 Introduction
-Large language models (LLMs) (Hadi et al.,, 2023; Brown et al.,, 2020; OpenAI,, 2023; Touvron et al., 2023a, ; Touvron et al., 2023b, ; Scao et al.,, 2022; Reid et al.,, 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks.
-A key principle, the Scaling Law (Kaplan et al.,, 2020), suggests that LLMs exhibit emergent abilities as model size increases, enhancing their capacity to understand context and handle long sequences (Xiong et al.,, 2023).
-This capacity growth allows LLMs to generate coherent and contextually accurate responses and enables various downstream applications, such as document summarization (Zhang et al.,, 2019; Zhang et al., 2024a, ), code generation (Chen et al., 2021b, ), and conversational AI (Bordes et al.,, 2016; OpenAI,, 2022), .
-Despite their success in various applications, the generation of LLMs incurs significant expenses, which escalate with increasing model size and sequence length.
-Notably, both the training (Strubell et al.,, 2020; Hoffmann et al.,, 2022; Dong et al., 2024a, ) and inference (Ainslie et al.,, 2023) stages involve frequent generation by LLMs, further contributing to these costs.
-Consequently, efficient LLMs have gained popularity in recent years (Hu et al.,, 2021; Wan et al.,, 2023).
-To address these challenges, quantization (Frantar et al.,, 2022; Lin et al.,, 2024; Dettmers et al.,, 2024; Xu et al.,, 2023) and pruning methods (Frankle and Carbin,, 2018; Blalock et al.,, 2020) are employed to reduce model size.
-Additionally, managing long sequences presents another cost due to the transformer attention mechanism.
-The quadratic complexity of the attention mechanism results in substantial computational burden when dealing with long sequences, scaling poorly with sequence length.
-Therefore, effective management of long sequences is essential for the practical deployment of LLMs.
-In this paper, we focus on the long-context scenario and aim to reduce the memory consumption associated with lengthy sequences.
-Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size of each head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference.
-To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or performing quantization over the caches.
+Large language models (LLMs) (Hadi et al., 2023; Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023a; b; Scao et al., 2022; Reid et al., 2024) have emerged as a dominant paradigm in natural language processing, achieving state-of-the-art performance across various tasks.
+A key principle, the Scaling Law (Kaplan et al., 2020), suggests that LLMs exhibit emergent abilities as model size increases, improving their capacity to understand complex context and handle long sequences (Xiong et al., 2023).
+This growth in capacity enables LLMs to generate coherent, contextually accurate responses and supports a variety of downstream applications, such as document summarization (Zhang et al., 2019; 2024a), code generation (Chen et al., 2021b), solving mathematical problems (Hendrycks et al., 2021; Zhou et al., 2023; Wang et al., 2023; Lightman et al., 2023), and conversational AI (OpenAI, 2022; 2023).
+Despite their success in various applications, generating outputs with LLMs incurs significant computational and financial costs, which rise with increasing model size and sequence length.
+Both the training (Strubell et al., 2020; Hoffmann et al., 2022; Dong et al., 2024a) and inference (Ainslie et al., 2023) stages involve frequent generation, further contributing to these costs.
+Consequently, efficient LLMs have gained traction in recent years (Hu et al., 2021; Wan et al., 2023).
+To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) are employed to reduce model size.
+Additionally, the key-value (KV) cache, stored in GPU memory alongside model parameters, scales linearly with both sequence length and batch size, creating a substantial memory burden when handling long sequences.
+Consequently, effective management of extended contexts is essential for the practical deployment of LLMs.
+In this paper, we focus on the long-context scenario, aiming to reduce memory consumption associated with processing lengthy sequences.
+Specifically, the number of KV cache parameters is the product of batch size $B$, sequence length $S$, number of layers $L$, number of heads $N$, channel size per head $D$, i.e., $\mathbf{K},\mathbf{V}\in\mathbb{R}^{B\times S\times L\times N\times D}$, which need to be stored in the GPU memory during inference.
+To reduce memory and computational costs during inference, efficiency can only be achieved by pruning the dimensions across $S,L,N,D$ or applying quantization to the caches.
 It is well-acknowledged that token importance tends to be sparse.
-Consequently, KV eviction algorithms have been proposed to minimize the KV cache memory cost from dimension $S$ (Xiao et al., 2023b, ; Li et al.,, 2024; Zhang et al., 2024c, ; Leviathan et al.,, 2023).
-Additionally, inter-layer redundancy has been explored (Liu et al., 2024a, ; Wu and Tu,, 2024; Brandon et al.,, 2024) to address the layer dimension $L$.
+Consequently, KV eviction algorithms have been proposed to reduce the memory footprint by addressing the sequence length dimension $S$ (Xiao et al., 2023b; Li et al., 2024; Zhang et al., 2024c; Leviathan et al., 2023).
+Additionally, inter-layer redundancy has been explored (Liu et al., 2024a; Wu & Tu, 2024; Brandon et al., 2024) to address the layer dimension $L$.
 Despite these advances, existing methods have largely overlooked the channel dimension $D$.
-In this paper, we highlight that the magnitude across Key cache channel dimensions is significantly unbalanced.
-Additionally, we observe that the attention weights exhibit a low-rank structure.
+In this paper, we highlight that the magnitudes across key cache channel dimensions are significantly imbalanced, and we observe a low-rank structure in attention weights.
 Based on these findings, we hypothesize that the channel dimension of the key cache exhibits redundancy.
-Consequently, we explore the redundancy of KV cache tensors specifically from dimension $D$, aiming to devise effective strategies that reduce costs without compromising performance.
-In our paper, we propose a simple yet effective method – ThinK, for KV cache pruning.
-To pinpoint the least significant channels, we formulate the task as an optimization problem, aiming to minimize the loss in attention weights attributable to pruning.
-To effectively address this problem, we establish a novel query-dependent criterion that assesses the importance of each channel.
+Consequently, we focus on exploring the redundancy in the KV cache along dimension $D$, aiming to develop strategies that reduce memory costs without compromising performance.
+In this paper, we introduce ThinK, a simple yet effective method for KV cache pruning.
+To pinpoint the least significant channels, we formulate the problem as an optimization task, aiming to minimize the loss in attention weights caused by pruning.
+To effectively address this problem, we propose a novel query-dependent criterion that assesses the importance of each channel.
 Using this criterion, we then select the most critical channels in a greedy fashion.
-We evaluate ThinK using the LLaMA3 (Meta,, 2024) and Mistral (Jiang et al.,, 2023) models and validate its effectiveness across various long-sequence datasets.
-The results indicate that when paired with token eviction methods, ThinK not only achieves comparable or superior accuracy but also reduces KV cache memory costs by more than 20%.
+We evaluate ThinK using the LLaMA (Meta, 2024) and Mistral (Jiang et al., 2023) models, and validate its effectiveness across various long-sequence datasets.
+The results indicate that, when paired with token eviction and KV cache quantization methods, ThinK not only maintains comparable or superior accuracy but also reduces KV cache memory costs by over 20%.
 Contributions.
-This work pioneers the investigation into the sparsity structure of the channels in the KV cache.
-Specifically, we discovered that the activated key cache is sparse given a specific query.
+This work pioneers the investigation into the sparsity within the channels of the KV cache.
+Specifically, we uncover that the activated key cache is sparse for a given query.
 This insight allows us to prune the key cache channels using a query-induced norm.
-Inspired by this phenomenon, we introduce a plug-and-play method ThinK, the first channel pruning method specifically designed for KV cache.
-ThinK reduces the dimensionality of the cache channel, leading to linear savings in both memory and computational requirements.
-As plug-and-play techniques, our proposed ThinK is orthogonal to other KV cache compression schemes (e.g.
+Building on this insight, we introduce ThinK, the first channel pruning method specifically designed for KV cache.
+ThinK reduces the dimensionality of the cache channels, leading to linear savings in memory usage.
+As a plug-and-play technique, ThinK is orthogonal to other KV cache compression schemes (e.g.
 KV cache eviction, quantization).
-Our extensive experiments demonstrate the remarkable efficiency of ThinK on LLaMA-3 and Mistral models.
-Furthermore, we explore the promising potential of extending ThinK to value cache pruning (ThinKV) in specific scenarios, highlighting the broad applicability of our method.
+Our extensive experiments demonstrate ThinK’s remarkable efficiency on the LLaMA and Mistral models.
+Moreover, we explore the potential extension of ThinK to value cache pruning (ThinKV), showcasing the broad applicability of our method.
 Figure 1: An illustration of the pruning procedure of ThinK.
-Within each head, scores are calculated for each channel, and only the top T channels out of D are selected for retention.
-A binary channel mask, along with the pruned keys, is subsequently stored in the cache memory.
+Within each attention head, scores are computed for each channel, and only the top $T$ channels out of $D$ are selected for retention.
+A binary channel mask, along with the pruned keys, is then stored in the cache memory.
 2 Observations
-We identify several key observations that motivate our approach to prune the channel of KV-Cache.
-Specifically, we visualize the magnitude of KV-cache and perform singular value decomposition of the attention in LLaMA3-8B model.
-We show the visualization for certain heads and layers for illustration purposes.
-Magnitude of KV-Cache Channel
-Figure 2 visualizes the absolute values of key and value cache across the tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis..
-Consistent with previous findings (Lin et al.,, 2024; Xiao et al., 2023a, ; Liu et al., 2024b, ), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns.
-For instance, the absolute value around $50^{th}$ channel of key cache has much larger magnitudes than other channels for all tokens in layer $14$ (Figure 2 (a)).
-The same observation applies to $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 2 (c)).
-Given such an observation, Liu et al., 2024b proposed to perform quantization over the channels of the key cache based on such observations.
-In addition to channel quantization, the observation also suggests that it is feasible to prune away certain key cache channels that contribute less to the attention mechanism.
-Moreover, channel quantization and pruning techniques are orthogonal such that they can be utilized concurrently to enhance model efficiency.
-(a) Key Cache ($14^{th}$ layer)
-(b) Value Cache ($14^{th}$ layer)
-(c) Key Cache ($20^{th}$ layer)
-(d) Value Cache ($20^{th}$ layer)
-Figure 2: Magnitude of key and value cache for LLaMA3-8B.
-The first head of layer 14 and layer 20 of LLaMA3-8B is selected to visualize the magnitude of the key and value caches.
-We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
-Singular Value Analysis
-We conducted singular value decomposition (SVD) (Demmel,, 1997) on the attention of the specified layers to investigate their principal components.
+We identify several key observations that motivate our approach to pruning the channels of the KV cache.
+Specifically, we visualize the magnitude of the KV cache and perform singular value decomposition (SVD) on the attention mechanism of the LLaMA model.
+Magnitudes of KV cache channels.
+Figure 4 (in Appendix A) visualizes the absolute values of the KV cache across tokens in each channel [^1] We use the visualization code from https://github.com/jy-yuan/KIVI/tree/main/vis..
+Consistent with previous findings (Lin et al., 2024; Xiao et al., 2023a; Liu et al., 2024b), we observe that only certain channels have significant magnitudes in the key cache, whereas the value cache lacks obvious patterns.
+For instance, in layer $14$ (Figure 4 (a)), the magnitudes in the key cache are substantially higher around the $50^{th}$ channel across all tokens.
+A similar pattern is observed in the $50^{th}$ and $150^{th}$ channels of the first head in layer $20$ (Figure 4 (c)).
+Given such an observation, Liu et al.
+(2024b) proposed to perform quantization over the channels of the key cache.
+Beyond quantization, our findings suggest that certain key cache channels with smaller contributions to the attention mechanism can be pruned.
+Moreover, channel quantization and pruning are orthogonal techniques, meaning they can be applied concurrently to further improve model efficiency.
+Singular value analysis.
+We conducted singular value decomposition (SVD) (Demmel, 1997) on the attention weights of the specified layers to investigate their principal components.
 The singular values derived from SVD capture the effective rank of the attention matrix, indicating how the information is distributed across different components.
-$\mathbf{U},\mathbf{\Sigma},\mathbf{V}$ $=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}$ (1)
-Figure 3 (a) illustrates the energy distribution of the singular values, plotted in log scale to enhance the visibility of differences among them.
+Figure 5 (a) (in Appendix A) illustrates the energy distribution of the singular values, plotted on a logarithmic scale to enhance visibility of the differences.
 Notably, only a few singular values exhibit high energy levels exceeding $0.01$ across all heads and layers, highlighting their relative significance.
-This observation aligns with previous findings (Bhojanapalli et al.,, 2021), where a small subset of singular values often captures most of the information in attention mechanisms.
+This observation aligns with previous findings (Bhojanapalli et al., 2021), where a small subset of singular values often captures most of the information in attention mechanisms.
 In addition, the rapid decay of the energy suggests that a low-rank approximation can effectively capture the essential information in the key cache.
-Figure 3 (b), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy.
-These findings suggest that the attention matrix is inherently low-rank (Wang et al.,, 2020; Chen et al., 2021a, ; Dong et al., 2024b, ), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al.,, 2024).
-(a) Energy of singular values
-(b) Cumulative energy of singular values
-Figure 3: The energy and cumulative energy of the singular values.
+Figure 5 (b) (in Appendix A), the normalized cumulative energy sum reveals that the top $50$ singular values account for over $90$% of the total energy.
+These findings suggest that the attention matrix is inherently low-rank (Wang et al., 2020; Chen et al., 2021a; Dong et al., 2024b), indicating that the key cache can be approximated using low-dimensional vectors (Singhania et al., 2024).
 3 ThinK
-Notations.
 […]
@@
 In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size.
-Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows:
-$2\times B\times S\times L\times N\times D,$ (2)
-where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension.
+Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows: $2\times B\times S\times L\times N\times D$, where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension.
 The KV cach size grows linearly as the batch size $B$ and sequence length $S$.
-For a model with multihead attention (MHA) (Vaswani et al.,, 2017), such as LLaMA2-7B (Touvron et al., 2023b, ), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters.
-The KV cache must be transferred from off-chip memory (HBM) (Jia et al.,, 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck.
+For a model with multihead attention (MHA) (Vaswani et al., 2017), such as LLaMA2-7B (Touvron et al., 2023b), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters.
+The KV cache must be transferred from off-chip memory (HBM) (Jia et al., 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck.
 This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies.
-Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b, ; Zhang et al., 2024c, ; Li et al.,, 2024) and precision (Hooper et al.,, 2024; Liu et al., 2024b, ).
+Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b; Zhang et al., 2024c; Li et al., 2024) and precision (Hooper et al., 2024; Liu et al., 2024b).
 We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
-Table 1: Performance comparison of pruning K cache by $l_{p}$ norm on LongBench.
+Table 1: Performance comparison of pruning key cache by $l_{p}$ norm on LongBench.
 Method $\lambda$ Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 RB-P
-H2O 0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
-+$l_{1}$ 30 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
-+$l_{1}$ 40 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
-+$l_{2}$ 30 23.98 17.04 35.19 39.27 31.29 20.4 21.62 22.46 23.34 40.5 89.75 40.71 5.54 68.67 60.12 58.52 37.40
-+$l_{2}$ 40 23.76 16.23 32.19 40.23 32.13 20.69 21.3 22.25 23.2 39.5 89.61 40.24 5.66 69.0 60.09 59.45 37.22
-SnapKV 0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
-+$l_{1}$ 30 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
-+$l_{1}$ 40 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
-+$l_{2}$ 30 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.5 90.33 40.31 5.7 68.42 62.65 60.07 40.28
-+$l_{2}$ 40 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
-Magnitude base Pruning: Based on the observations in Figure 2 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
+H2O 0.0 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
++$l_{1}$ 0.3 23.38 17.15 34.99 40.56 31.49 19.90 21.37 22.13 23.44 40.50 90.10 40.65 5.41 69.00 58.64 54.99 37.11
++$l_{1}$ 0.4 23.51 15.40 34.37 40.71 31.28 20.24 21.25 22.29 22.54 38.50 89.22 39.27 5.87 68.33 58.47 54.33 36.60
++$l_{2}$ 0.3 23.98 17.04 35.19 39.27 31.29 20.40 21.62 22.46 23.34 40.50 89.75 40.71 5.54 68.67 60.12 58.52 37.40
++$l_{2}$ 0.4 23.76 16.23 32.19 40.23 32.13 20.69 21.30 22.25 23.20 39.50 89.61 40.24 5.66 69.00 60.09 59.45 37.22
+SnapKV 0.0 24.84 23.96 38.77 42.75 34.55 20.87 22.26 22.61 23.97 70.00 90.52 40.29 5.81 69.50 59.04 51.81 40.10
++$l_{1}$ 0.3 24.43 24.63 40.11 41.83 33.47 21.22 21.47 22.41 23.73 66.50 90.39 40.20 5.70 68.10 61.04 55.37 40.04
++$l_{1}$ 0.4 24.58 24.87 39.30 42.76 31.95 20.47 20.95 22.22 23.42 55.50 90.22 39.13 5.82 68.39 60.71 56.10 39.15
++$l_{2}$ 0.3 24.47 24.73 38.16 41.86 32.23 20.23 21.59 22.45 23.77 67.50 90.33 40.31 5.70 68.42 62.65 60.07 40.28
++$l_{2}$ 0.4 24.52 23.75 38.35 42.42 32.96 20.39 21.21 22.28 23.41 60.00 90.20 39.59 5.75 68.29 61.96 60.59 39.74
+Magnitude based Pruning: Based on the observations in Figure 4 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
 $M_{n,d}=\big{\|}K[n,:,d]\big{\|}_{p}.$ (1)
-Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head:
-$I=\textbf{Top}_{T}(M,T)$ (4)
-where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel.
+Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head: $I=\textbf{Top}_{T}(M,T)$ where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel.
 $n\in[1,N]$ and $d\in[1,D]$ are indicators of heads and channels in key cache.
 $I\in\mathbb{(Z^{+})}^{N\times T}$ stores the indicators of the top $T$ values in tensor $M$ per head.
-In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA3-8B model.
-We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al.,, 2023).
-Compared to the baseline methods, H2O (Zhang et al., 2024c, ) and SnapKV (Li et al.,, 2024), both with a KV length of 512, we further prune the channels of the key cache.
+In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA-3-8B model.
+We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al., 2023).
+Compared to the baseline methods, H2O (Zhang et al., 2024c) and SnapKV (Li et al., 2024), both with a KV length of 512, we further prune the channels of the key cache.
 A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning.
@@
 3.2 Query-Driven Pruning
-For each head, the attention scores are computed using the queries and keys, and then applied to the values. The formula for the attention for head $i$ is:
-$\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}$ (5)
-where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$.
-When pruning one channels of $\textbf{K}_{i}$, the corresponding channels in $\textbf{Q}_{i}$ will also be removed.
+For each head, the attention scores are computed using the queries and keys, and then applied to the values.
+The formula for the attention for head $i$ is: $\text{Attention}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})=\text{softmax}(\frac{\textbf{Q}_{i}\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i},$ where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$.
+When one channel of $\textbf{K}_{i}$ is pruned, the corresponding channel in $\textbf{Q}_{i}$ will also be removed.
 We aim to find the optimal subset of channels to prune, denoted by the selection matrix $\textbf{S}\in\{0,1\}^{D\times D}$, where S is a diagonal matrix with binary entries ($1$ for keeping a channel, 00 for pruning it).
-To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights:
-$\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}$ (6)
-Given a pruning ratio $\lambda$, it can further expanded as:
-$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$
+To better maintain the performance after pruning the channels, we minimize the Frobenius norm of the difference between the original and pruned attention weights: $\min_{\textbf{S}}\|\textbf{Q}_{i}\textbf{K}^{T}_{i}-\textbf{Q}_{i}\textbf{S}(\textbf{K}_{i}\textbf{S})^{T}\|_{F}.$ Given a pruning ratio $\lambda$, it can further expanded as:
+$\min_{\mathbf{S}}$ $\quad\left\|\mathbf{Q}_{i}\mathbf{K}_{i}^{T}-\mathbf{Q}_{i}\mathbf{S}\mathbf{K}_{i}^{T}\right\|_{F}$ (2)
 subject to $\quad\text{trace}(\mathbf{S})=\lfloor(1-\lambda)D\rfloor$
-$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$ (7)
+$\quad\mathbf{S}=\text{diag}(s_{1},s_{2},\ldots,s_{D}),\text{ where }s_{j}\in\{0,1\}$
 For simplicity, we use greedy algorithm to optimize $\mathbf{S}$.
-To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores:
-$\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F}$ (8)
-$I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$ (9)
+To achieve the pruning goal, we define a criterion for evaluating the importance of each channel and greedily select the channels with largest scores: $\text{Score}_{i}[j]=\big{\|}\textbf{Q}_{i}[:,j]\textbf{K}_{i}[:,j]^{T}\big{\|}_{F},\quad I_{i}=\textbf{Top}_{T}(\text{Score}_{i},T)$.
 Here’s a detailed explanation of why it optimizes the selection matrix.
@@
 Observation Window.
-Based on the observations in SnapKV (Li et al.,, 2024) that the last window of input sequence recognizes highly similar attention allocation pattern with the actual generation.
+Based on the observations in SnapKV (Li et al., 2024) that the last window of input sequence recognizes highly similar attention pattern with generation.
 To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
+3.3 Implementation of ThinK
 Figure 2: Implementation during decoding.
-3.3 Implementations
 Following SnapKV, we focus on the long context input scenario.
@@
 Additionally, we maintain a binary mask to indicate which channels have been pruned.
-Note that the memory overhead associated with this mask is negligible.
-Figure 4 illustrates one implementation of our method during the decoding stage.
-The pruned keys are first zero-filled using the mask to restore them to their original size, and then concatenated with the unpruned keys.
-This implementation can be effectively integrated with optimization techniques such as FlashAttention (Dao,, 2023).
-The second implementation involves initially pruning the Query using the mask.
+it is important to note that the memory overhead associated with this mask is negligible.
+Figure 2 illustrates one implementation of ThinK during the decoding stage.
+This implementation begins by pruning the query using the mask.
 The pruned query is then multiplied by the pruned key, while the unpruned query is applied to the unpruned Key.
 Subsequently, the two outputs are concatenated.
-In theory, this implementation has lower computational costs.
-4 Experiment Results
-In this section, we conduct extensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction.
-First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1.
-Next, we report the performance of our KV pruning method with two different KV compression methods in long-context scenarios in Section 4.2.
-Finally, we present the results of ablation experiments in Section 4.4.
-4.1 Settings
-Benchmark Datasets.
-We perform evaluations of our method with state-of-the-art KV cache compression methods on two widely used benchmarks: LongBench and Needle-in-a-Haystack.
-LongBench (Bai et al.,, 2023) is designed to comprehensively evaluate LLM’s long context understanding capabilities.
-It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion.
-The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference.
-Needle-in-a-Haystack (Kamradt,, 2023) is a recent popular test challenge that requires models to accurately identify a small piece of information (“needle”) in a long document (“haystack”), where the needle is placed at a random position.
-This challenge can test if KV cache compression methods still retain the small piece of critical information.
-Baseline Approaches.
-The baseline methods include Heavy Hitter Oracle (H2O) and SnapKV, both of which are the state-of-the-art KV cache compression methods but use different strategies for selecting KV cache.
-H2O (Zhang et al., 2024c, ) is designed to reduce memory usage by dynamically balancing recent tokens and Heavy Hitter (H2) tokens, where H2 tokens are a small set of tokens that contribute most of the value when computing attention scores.
-SnapKV (Li et al.,, 2024) automatically compresses KV caches by selecting clustered important KV positions for each attention head.
-This is inspired by the fact that each attention head has a specific pattern, focusing on certain attention features during generation.
-Table 2: Performance comparison of Key cache pruning on LLaMA3-8B by ThinK on LongBench.
+Table 2: Performance comparison of key cache pruning on LLaMA-3-(8B/70B)-Instruct on LongBench. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$.
 Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 RB-P
-KV-size Full
+LLaMA-3-8B-Instruct, KV-size Full
 ALL KV 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
-KV-size 128
+LLaMA-3-8B-Instruct, KV-size 128
 H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
-+ThinK(40) 22.85 14.55 29.49 38.63 30.84 18.90 20.12 21.96 20.68 38.50 86.38 38.40 5.50 69.17 57.93 56.12 35.63
-+ThinK(50) 23.47 14.06 28.67 38.35 30.21 17.87 19.69 21.94 19.95 38.50 87.14 38.07 4.92 69.50 57.99 56.66 35.44
++ThinK (0.4) 22.85 14.55 29.49 38.63 30.84 18.90 20.12 21.96 20.68 38.50 86.38 38.40 5.50 69.17 57.93 56.12 35.63
++ThinK (0.5) 23.47 14.06 28.67 38.35 30.21 17.87 19.69 21.94 19.95 38.50 87.14 38.07 4.92 69.50 57.99 56.66 35.44
 […]
-+ThinK(40) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
-+ThinK(50) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
-KV-size 512
++ThinK (0.4) 22.11 14.67 32.49 36.25 28.63 18.80 18.93 21.49 20.14 44.50 88.11 38.32 5.75 69.17 58.21 55.89 35.84
++ThinK (0.5) 21.79 14.73 32.03 37.52 27.86 18.28 18.50 21.52 19.71 43.50 86.00 38.35 5.59 69.50 57.96 56.96 35.61
+LLaMA-3-8B-Instruct, KV-size 512
 H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
-+ThinK(40) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
-+ThinK(50) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
-+ThinK(60) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
++ThinK (0.4) 23.76 17.80 33.80 40.39 30.70 19.09 21.82 22.51 23.78 41.00 90.16 40.67 5.15 69.25 60.77 57.58 37.39
++ThinK (0.5) 24.17 16.96 35.76 39.47 30.29 18.67 21.39 22.59 23.06 41.00 89.81 40.35 5.23 69.33 60.20 58.34 37.29
++ThinK (0.6) 23.40 14.83 32.62 38.47 30.97 19.81 20.80 22.04 21.60 40.00 88.79 38.90 5.36 69.50 58.28 57.65 36.44
 […]
-+ThinK(40) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
-+ThinK(50) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
-+ThinK(60) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
-KV-size 1024
++ThinK (0.4) 24.58 25.44 37.03 41.87 33.45 20.58 21.77 22.42 24.16 70.00 90.39 40.29 6.06 69.50 62.05 59.23 40.55
++ThinK (0.5) 24.85 25.10 37.06 41.58 32.34 20.60 21.61 22.44 23.66 69.50 90.39 39.70 5.84 69.79 61.57 59.42 40.34
++ThinK (0.6) 25.98 22.77 38.37 40.44 33.19 19.90 20.84 22.21 22.55 59.00 90.32 38.12 6.39 69.50 59.14 58.40 39.20
+LLaMA-3-8B-Instruct, KV-size 1024
 H2O 25.62 22.16 36.81 41.01 33.53 19.41 23.28 22.65 25.41 46.50 90.82 41.78 5.79 69.25 59.69 55.50 38.70
-+ThinK(40) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
-+ThinK(50) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
-+ThinK(60) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
++ThinK (0.4) 25.52 21.93 37.17 41.56 31.22 20.17 22.89 22.95 25.21 47.00 90.74 41.34 5.57 69.50 62.58 58.67 39.00
++ThinK (0.5) 25.41 22.19 37.64 40.92 31.27 18.66 22.17 22.22 24.84 46.50 90.34 40.59 5.20 69.50 61.71 57.99 38.57
++ThinK (0.6) 24.06 17.80 37.85 38.63 29.98 19.40 21.41 22.32 23.42 44.50 90.16 39.43 5.84 69.50 58.31 58.73 37.58
 […]
-+ThinK(40) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
-+ThinK(50) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
-+ThinK(60) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
-KV-size 2048
++ThinK (0.4) 24.88 27.72 38.60 43.16 32.44 20.67 24.21 22.79 25.56 71.50 90.45 40.94 5.93 69.50 62.77 59.45 41.29
++ThinK (0.5) 24.82 27.26 39.66 42.82 32.09 19.56 23.52 22.48 25.34 71.50 90.43 40.74 5.20 69.50 62.46 59.75 41.07
++ThinK (0.6) 24.46 27.35 38.22 41.96 31.64 20.18 21.89 22.83 23.68 70.00 90.19 38.69 6.10 69.50 58.87 59.26 40.30
+LLaMA-3-8B-Instruct, KV-size 2048
 H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
-+ThinK(40) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
-+ThinK(50) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
-+ThinK(60) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
++ThinK (0.4) 25.56 26.31 39.20 42.96 31.81 20.53 24.23 23.35 25.90 53.50 90.56 41.03 5.52 69.25 62.10 59.00 40.05
++ThinK (0.5) 25.01 25.37 38.82 42.32 31.27 20.50 23.78 23.21 26.03 53.00 90.37 40.86 5.13 69.50 61.91 58.95 39.75
++ThinK (0.6) 24.37 22.14 37.77 40.13 29.50 20.26 22.09 22.76 24.78 49.50 90.16 39.69 5.56 69.50 59.24 58.78 38.51
 […]
-+ThinK(40) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
-+ThinK(50) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
-+ThinK(60) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
-Implementation Details.
-In this paper, we use LLaMA3-8B-Instruct (Meta,, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al.,, 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al.,, 2020).
-Our ThinK aims to prune channels of the KV, which is agnostic to KV cache compression methods.
-If there is no other statement, we prune the key cache by default.
-All the experiments are conducted using one Nvidia A100.
-To fairly compare KV cache compression methods and KV cache compression integrated with ThinK, we use the same hyperparameters for both.
-For example, when comparing SnapKV and SnapKV integrated with ThinK, we set the maximum pooling kernel size to 7 and the observation window size to 32, using the same KV-size for both.
-Table 3: Performance comparison of Key cache pruning on Mistral-7B by ThinK on LongBench.
++ThinK (0.4) 25.41 29.79 39.21 43.35 33.96 21.49 25.78 23.11 26.23 73.00 90.56 41.79 5.81 69.50 62.45 59.19 41.91
++ThinK (0.5) 25.00 30.25 39.27 43.23 32.93 21.24 25.16 23.01 26.5 73.00 90.37 41.26 5.45 69.50 62.3 59.84 41.77
++ThinK (0.6) 24.89 28.88 40.44 41.30 29.99 21.34 23.48 22.9 24.99 72.50 90.36 38.5 5.71 69.50 59.77 59.50 40.88
+LLaMA-3-70B-Instruct, KV-size 128
+SnapKV 25.91 39.41 43.83 49.60 51.23 27.76 22.14 21.91 23.16 69.00 91.55 43.54 12.50 72.00 48.41 63.49 44.09
++ThinK (0.4) 25.64 39.20 43.60 50.22 50.50 29.32 21.70 21.96 23.35 68.00 91.27 43.24 12.50 73.00 48.01 62.43 44.00
++ThinK (0.5) 26.31 38.76 44.86 48.54 49.62 28.97 21.46 22.01 22.91 67.00 91.52 43.15 12.50 72.50 47.21 60.82 43.63
+LLaMA-3-70B-Instruct, KV-size 512
+SnapKV 27.95 45.19 48.50 50.97 54.53 29.78 25.34 22.36 26.03 73.50 92.63 45.07 12.50 72.50 45.21 68.22 46.27
++ThinK (0.4) 27.47 45.31 48.57 51.22 54.32 30.05 25.42 22.72 26.20 73.50 92.13 45.53 12.50 73.00 48.32 66.99 46.45
++ThinK (0.5) 26.97 44.55 48.16 50.84 53.80 30.57 25.29 22.65 25.53 73.00 92.13 43.66 12.50 73.00 50.52 64.82 46.12
+LLaMA-3-70B-Instruct, KV-size 1024
+SnapKV 26.80 46.21 49.93 51.70 54.71 29.86 27.61 22.43 27.15 73.50 92.38 46.18 12.50 72.50 42.84 69.89 46.64
++ThinK (0.4) 27.04 46.01 50.13 51.96 54.36 29.87 27.74 22.78 27.07 73.50 91.88 46.35 12.50 73.00 45.05 67.87 46.69
++ThinK (0.5) 27.62 46.22 48.97 51.79 53.39 30.47 27.45 23.05 26.57 73.50 91.88 43.99 12.50 72.50 47.41 66.84 46.51
+LLaMA-3-70B-Instruct, KV-size 2048
+SnapKV 27.44 46.51 49.60 51.80 54.77 31.05 29.67 22.44 27.43 73.50 92.38 45.98 12.50 72.50 41.86 68.72 46.76
++ThinK (0.4) 27.13 46.26 50.04 51.72 55.03 31.19 29.75 22.47 27.28 73.50 91.88 46.37 12.50 72.50 42.66 67.77 46.75
++ThinK (0.5) 27.84 46.86 49.18 51.97 53.58 31.44 29.41 22.89 27.33 73.50 91.88 43.60 12.50 72.50 44.78 66.65 46.62
+Table 3: Performance comparison of key cache pruning on Mistral-7B-Instruct-v0.2 on LongBench. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$.
 Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 RB-P
-KV-size Full
+Mistral-7B-Instruct-v0.2, KV-size Full
 ALL KV 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
-KV-size 128
+Mistral-7B-Instruct-v0.2, KV-size 128
 H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
-+ThinK(40) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
-+ThinK(50) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
-+ThinK(60) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
++ThinK (0.4) 21.17 21.90 39.29 29.92 20.99 12.30 20.84 22.91 21.92 39.00 82.70 40.35 2.97 79.21 51.19 48.32 34.69
++ThinK (0.5) 21.67 21.80 39.48 28.74 20.65 13.14 20.57 22.83 21.78 39.00 82.54 40.12 3.61 78.39 50.27 48.4 34.56
++ThinK (0.6) 21.04 21.30 39.56 28.68 21.29 13.97 20.13 22.52 21.81 39.50 82.05 39.14 4.16 74.23 49.83 47.67 34.18
 […]
-+ThinK(40) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
-+ThinK(50) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
-+ThinK(60) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
-KV-size 512
++ThinK (0.4) 20.52 21.00 42.65 37.58 22.09 15.23 19.29 22.01 21.22 47.00 83.85 39.64 3.20 67.45 51.48 48.31 35.16
++ThinK (0.5) 20.67 20.60 43.37 37.27 21.58 15.66 19.06 21.79 21.02 47.00 83.38 39.77 3.65 67.06 50.80 48.35 35.06
++ThinK (0.6) 21.25 20.82 44.20 36.21 21.68 16.47 19.05 21.99 20.73 45.00 83.81 38.79 4.19 66.90 49.99 47.61 34.92
+Mistral-7B-Instruct-v0.2, KV-size 512
 H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
-+ThinK(40) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
-+ThinK(50) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
-+ThinK(60) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
++ThinK (0.4) 21.58 26.15 44.49 32.73 23.99 15.09 23.56 23.28 24.45 42.00 85.58 42.58 3.18 85.7 54.39 51.15 37.49
++ThinK (0.5) 22.76 25.74 44.61 31.74 23.25 13.91 23.31 23.13 24.34 41.00 85.39 41.85 2.82 84.36 54.69 50.88 37.11
++ThinK (0.6) 22.91 25.57 44.04 29.48 22.88 13.67 23.31 22.64 24.10 41.00 85.31 41.15 2.98 82.34 53.70 50.25 36.58
 […]
-+ThinK(40) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
-+ThinK(50) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
-+ThinK(60) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
-KV-size 1024
++ThinK (0.4) 24.27 28.46 49.26 38.13 24.22 16.92 23.59 23.70 24.46 67.50 85.9 42.51 2.92 85.32 55.89 53.35 40.40
++ThinK (0.5) 24.56 29.22 48.59 37.70 24.27 17.39 23.68 23.65 24.58 67.50 86.05 42.01 3.07 86.30 56.49 53.29 40.52
++ThinK (0.6) 24.07 28.27 49.10 38.65 24.31 17.52 23.16 23.51 24.23 67.00 86.33 40.78 3.69 83.74 54.94 52.23 40.10
+Mistral-7B-Instruct-v0.2, KV-size 1024
 H2O 23.67 28.55 46.40 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
-+ThinK(40) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
-+ThinK(50) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
-+ThinK(60) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
++ThinK (0.4) 23.97 28.91 45.84 35.78 24.88 14.55 25.11 23.35 25.83 45.50 86.11 42.44 3.23 84.82 56.21 53.02 38.72
++ThinK (0.5) 23.89 28.40 46.60 35.57 24.26 14.78 24.98 23.31 25.68 44.50 86.16 42.72 3.38 83.20 55.88 52.63 38.50
++ThinK (0.6) 23.87 27.76 46.25 35.28 24.38 14.74 24.35 23.35 25.50 44.50 85.38 41.37 3.34 81.42 55.21 51.89 38.04
 […]
-+ThinK(40) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
-+ThinK(50) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
-+ThinK(60) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
-KV-size 2048
++ThinK (0.4) 25.22 30.48 48.58 41.11 25.28 18.99 25.91 24.00 26.13 70.00 86.64 43.35 2.98 86.3 56.71 54.19 41.62
++ThinK (0.5) 25.63 30.08 49.41 40.59 25.13 19.58 25.47 24.23 25.92 69.5 86.67 42.31 2.74 84.78 57.43 53.59 41.44
++ThinK (0.6) 24.69 29.3 48.90 40.44 25.33 19.58 25.23 23.6 25.25 69.00 86.85 40.86 3.19 83.70 56.3 53.30 40.97
+Mistral-7B-Instruct-v0.2, KV-size 2048
 H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.00 86.35 42.48 2.72 86.64 56.98 53.91 40.69
-+ThinK(40) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
-+ThinK(50) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
-+ThinK(60) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
++ThinK (0.4) 25.40 30.80 48.45 39.64 26.08 16.82 27.12 23.79 26.65 53.50 86.39 43.03 3.29 86.39 56.61 53.60 40.47
++ThinK (0.5) 25.68 31.24 48.69 39.65 25.84 16.72 26.69 23.57 26.78 52.00 86.74 42.85 4.01 83.46 57.12 53.67 40.29
++ThinK (0.6) 25.83 31.00 48.23 38.58 25.71 16.54 26.51 23.81 26.28 50.50 86.57 42.05 3.36 82.49 56.04 52.67 39.76
 […]
-+ThinK(40) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
-+ThinK(50) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
-+ThinK(60) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52
++ThinK (0.4) 25.77 32.67 48.70 41.06 27.07 19.14 28.91 24.37 26.88 70.00 86.37 42.75 3.61 87.38 57.21 54.44 42.27
++ThinK (0.5) 26.44 32.94 49.02 40.86 26.84 19.49 28.46 24.51 26.72 70.00 86.50 41.75 2.78 84.70 56.47 54.15 41.98
++ThinK (0.6) 26.00 32.53 48.73 40.95 26.77 18.92 27.40 23.97 26.37 70.00 86.45 41.12 3.31 82.24 56.01 53.53 41.52
+4 Experiment Results
+In this section, we conduct comprehensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction.
+First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1.
+Next, we report the performance ThinK with two KV eviction methods and one KV quantization methods in long-context scenarios in Section 4.2 and  4.3.
+Finally, we present the results of ablation experiments in Section 4.4.
+4.1 Settings
+Benchmark Datasets.
+We evaluate our proposed method against state-of-the-art KV cache compression methods on two widely recognized benchmarks: LongBench and Needle-in-a-Haystack.
+LongBench (Bai et al., 2023) is designed to comprehensively assess the long context understanding capabilities of LLMs.
+It includes 17 datasets covering six different tasks: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion.
+The average input length of LongBench is 6,711 words, which necessitates reducing the KV cache to lower memory usage for inference.
+Needle-in-a-Haystack (Kamradt, 2023) is a recently developed benchmark that tests a model’s ability to accurately locate a small but crucial piece of information (the ”needle”) embedded within a lengthy document (the ”haystack”).
+The random positioning of the needle in this challenge serves as a critical test to determine whether KV cache compression methods can retain essential information without loss of accuracy.
+Baseline Approaches.
+The baseline methods in our evaluations include Heavy Hitter Oracle (H2O), SnapKV and KIVI, all of which are the state-of-the-art KV cache compression methods but use different strategies.
+H2O (Zhang et al., 2024c) is designed to reduce memory usage by dynamically managing the balance between recent tokens and Heavy Hitter (H2) tokens.
+H2 tokens represent a small set of tokens that contribute most of the value when computing attention scores.
+SnapKV (Li et al., 2024) introduces an automated compression mechanism that selects clustered, important KV positions for each attention head, optimizing the KV cache without sacrificing performance.
+KIVI (Liu et al., 2024b) reduces memory overhead by quantizing the KV cache into lower-precision formats, significantly lowering the memory cost while preserving model accuracy.
+Implementation Details.
+In this paper, we use LLaMA-2-7B-chat, LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct (Meta, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al., 2020).
+Our ThinK aims to prune channels of the key cache, which is agnostic to KV cache compression methods.
+If there is no other statement, we prune the key cache by default.
+All the experiments are conducted using NVIDIA A100 GPUs.
+To ensure a fair comparison between KV cache compression methods and their integration with ThinK, we applied consistent hyperparameters across both settings.
+For instance, when comparing SnapKV and SnapKV integrated with ThinK, we used a maximum pooling kernel size of 7 and an observation window size of 32, maintaining the same KV-size for both configurations.
+Table 4: Performance evaluation of combining ThinK with KIVI Liu et al. (2024b) on LongBench. ThinK ($0.4$) indicates we prune the key cache channels with a pruning ratio of $\lambda=0.4$.
+Method   Bit Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
+NrtvQA
+Qasper
+MF-en
+HotpotQA
+2WikiMQA
+Musique
+GovReport
+QMSum
+MultiNews
+TREC
+TriviaQA
+SAMSum
+PCount
+PRe
+Lcc
+RB-P
+KIVI 2/2 19.47 18.62 30.28 29.42 25.00 10.30 21.34 20.51 25.10 63.00 85.04 40.16 4.00 8.00 58.04 52.48 31.92
++ThinK (0.4) 2/2 19.46 19.01 30.52 28.79 25.78 9.53 22.11 20.66 25.73 63.00 84.62 41.54 3.50 7.00 56.51 48.92 31.77
 4.2 Results on LongBench
-Tables 2 and  3 present the results of KV compression methods and KV compression methods integrated with our channel pruning method of Key cache (ThinK) over two different base LLMs across various KV-sizes on LongBench.
-The following observations can be drawn: (1) Our method can further prune the channels of the key cache after compressing the KV cache with H2O and SnapKV.
-For the base model LLaMA3-8B, our method reduces memory usage and slightly improves performance for both H2O and SnapKV.
-For the base model Mistral-7B, our method reduces memory with only a slight drop in performance in some cases.
-(2) Comparing the results of SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm shown in Table 1, our proposed query-based channel pruning shows better performance when the pruning ratio is $40\%$ compared to no pruning.
-However, $l_{1}$ and $l_{2}$ norm methods cannot outperform the KV compression methods without pruning.
-(3) When the pruning ratio is small, performance tends to be better than when the pruning ratio is large.
-(4) When the KV-size is increased from $128$ to $2048$, the performance of our channel pruning method improves.
-Notably, with a KV cache size of $2048$ and a pruning ratio of $40\%$, our method can even outperform the LLaMA3-8B with a full KV cache.
-The above observations indicate that our method is agnostic to existing KV cache compression methods and can further improve their performance and memory reduction.
-Additionally, query-based pruning is more effective than $l_{1}$ and $l_{2}$ norm for channel pruning in LLMs.
+Tables 2 and  3 present the results of KV compression methods and their integration with our proposed channel pruning technique for the key cache (ThinK) across three different base LLMs, evaluated at various KV-sizes on the LongBench benchmark.
+The following observations can be drawn: (1) Our method successfully prunes the channels of the key cache after the KV cache has been compressed using H2O and SnapKV.
+For the LLaMA-3-8B-Instruct base model, our approach reduces memory usage while slightly improving performance for both H2O and SnapKV.
+For the Mistral-7B-Instruct-v0.2 base model, our method similarly reduces memory usage, with only a minor performance drop in some cases.
+For the larger LLaMA-3-70B base model, our method achieves comparable or superior performance after pruning $40\%$ of the key cache channels, compared to the SnapKV baselines.
+(2) When Comparing SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm in Table 1, our query-driven channel pruning approach demonstrates superior performance when the pruning ratio of $\lambda=0.4$.
+(3) Lower pruning ratios generally result in better performance compared to higher pruning ratios.
+(4) As the KV-size increases from $128$ to $2048$, the performance of our channel pruning method improves.
+Notably, with a KV-size of $2048$ and a pruning ratio of $0.4$, our method even surpasses the performance of LLaMA-3-8B-Instruct with a full KV cache.
+These findings suggest that our method is agnostic to the underlying KV cache compression techniques and can further enhance both performance and memory efficiency.
+Moreover, query-driven channel pruning proves more effective than $l_{1}$ and $l_{2}$ norm-based methods for channel pruning in LLMs.
+We further validate the efficacy of our method by applying it to the KV cache quantization technique KIVI (Liu et al., 2024b), as shown in Table 4.
+First, we prune $40\%$ of the key cache channels, followed by quantization of the remaining channels into 2-bit (implementation details provided in Appendix C.1).
+Compared to the standard KIVI method, our approach reduces KV cache memory by $20\%$, with minimal performance degradation.
 4.3 Results on Needle-in-a-Haystack
-Table 4 presents the results of the Needle-in-a-Haystack test based on the SnapKV (Li et al.,, 2024) approach with KV cache size of input prompt from $128$ to $2048$.
-With a modest pruning ratio $\lambda=40$%, the accuracy of ThinK is consistently better or comparable than the original SnapKV across both LLaMA3 and Mistral with different KV cache sizes.
-Such comparisons demonstrate that the proposed query-driven channel pruning can effectively retain the informative channels and discard the noisy ones.
-With a pruning ratio of $\lambda\geq 50$%, we observe an accuracy drop with a small KV-cache size, particularly for $128$ and $512$ across both LLaMA3 and Mistral.
-However, ThinK is still able to obtain comparable performance with SnapKV when we have a larger KV cache size (i.e., $1024$ and $2048$).
-Intuitively, a larger pruning ratio with a smaller KV cache size could tend to lose more important information compared to a larger KV cache size.
-In addition, the performance on larger KV cache size also suggests ThinK is robust for long context retrieval tasks.
-Figure 6 (a)-(d) visualize the Needle-in-a-Haystack test accuracy against different token lengths and depth percent.
-The KV cache size is configured to $128$ and $1024$, with pruning ratio $\lambda$ equal to $40$% and $50$%, respectively.
-First, ThinK can maintain the original retrieval capability of SnapKV though the overall numerical accuracy is different (i.e., $77.8$ vs $78.6$ and $90.4$ vs $90.6$).
-Because ThinKachieves the same accuracy at most of the entries with different token limits and depths as Snap KV.
-In addition, ThinK also successfully retrieves some “needles” where SnapKV fails, leading to higher overall accuracy.
-The visualization results further demonstrate the robustness of ThinK in a fine-grained perspective and ThinK can enhance the original quantization approach.
 Table 5: Needles-in-a-Haystack Test Results
@@
 128 512 1024 2048
-LLaMA3-8B SnapKV 00 79.6 90.2 91.2 91.7
-SnapKV+ThinK 40 79.6 90.3 91.2 91.7
-SnapKV+ThinK 50 77.4 89.6 91.0 91.7
-Mistral-7B SnapKV 00 77.8 89.5 90.4 90.8
-SnapKV+ThinK 40 78.6 90.1 90.6 90.9
-SnapKV+ThinK 50 78.1 90.1 90.8 91.1
-SnapKV+ThinK 60 75.9 89.2 90.6 91.1
+LLaMA3-8B-Instruct SnapKV 0.0 79.6 90.2 91.2 91.7
+SnapKV+ThinK 0.4 79.6 90.3 91.2 91.7
+SnapKV+ThinK 0.5 77.4 89.6 91.0 91.7
+Mistral-7B-Instruct-v0.2 SnapKV 0.0 77.8 89.5 90.4 90.8
+SnapKV+ThinK 0.4 78.6 90.1 90.6 90.9
+SnapKV+ThinK 0.5 78.1 90.1 90.8 91.1
+SnapKV+ThinK 0.6 75.9 89.2 90.6 91.1
+Table 5 presents the results of the Needle-in-a-Haystack test, using the SnapKV (Li et al., 2024) approach with varying KV-sizes, ranging from $128$ to $2048$.
+With a modest pruning ratio of $\lambda=0.4$, ThinK consistently outperforms or matches the accuracy of the original SnapKV across both LLaMA-3 and Mistral models, regardless of KV-size.
+These comparisons demonstrate that the proposed query-driven channel pruning method effectively retains informative channels while discarding noisy ones.
+However, when the pruning ratio increases to $\lambda\geq 0.5$, we observe a drop in accuracy with smaller KV-sizes, particularly for $128$ and $512$, across both LLaMA-3 and Mistral models.
+Despite this, ThinK achieves comparable performance with SnapKV when the KV-size is larger(i.e., $1024$ and $2048$).
+Intuitively, a larger pruning ratio with a smaller KV-size may lead to the loss of more critical information compared to scenarios with a larger KV-size.
+In addition, the performance on larger KV-sizes suggests that ThinK is robust for long-context tasks.
+Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depth percentages.
+The KV-sizes are set to $128$ and $1024$, with pruning ratios of $\lambda=0.4$ and $\lambda=0.5$, respectively.
+ThinK preserves the retrieval capabilities of SnapKV, although there are minor numerical differences in overall accuracy (e.g., $77.8$ vs.
+$78.6$ and $90.4$ vs.
+$90.6$).
+ThinK matches SnapKV in accuracy for the majority of token limits and depths, demonstrating consistency in performance.
+Furthermore, ThinK successfully retrieves certain ”needles” that SnapKV fails to capture, resulting in improved overall accuracy.
+These visualizations highlight the robustness of ThinK from a fine-grained perspective, illustrating its capacity to enhance the original approach.
 4.4 Ablation Studies
-Table 5: Performance comparison of pruning K cache using different recent-size.
+Table 6: Performance comparison of key cache pruning with varying recent-sizes.
 Recent-Size Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 RB-P
-H2O + ThinK($\lambda=40\%$)
+H2O + ThinK ($\lambda=0.4$)
 0 24.40 27.50 45.42 35.17 24.45 13.02 27.65 23.88 26.86 53.50 86.06 41.73 3.01 83.42 55.12 51.32 38.91
@@
 128 25.69 30.93 48.32 39.63 26.08 16.82 27.18 23.92 26.62 53.50 86.39 42.96 3.29 86.39 56.77 53.60 40.51
-SnapKV + ThinK($\lambda=40\%$)
+SnapKV + ThinK ($\lambda=0.4$)
 0 24.94 28.58 45.78 39.59 25.40 15.92 29.50 24.05 26.72 70.00 85.60 41.38 2.97 84.00 55.27 52.39 40.76
@@
 128 25.75 32.49 48.61 41.01 27.18 19.14 28.79 24.64 26.77 70.00 86.37 42.77 3.61 87.13 57.19 54.39 42.24
-Table 6: ablations with the same memory consumption.
+Table 7: Performance comparison of key cache pruning with the same memory consumption.
 Methods Memory(M) Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 Impact of Different Recent Sizes.
-Preserve the most recent KV embeddings (Zhang et al., 2024c, ; Li et al.,, 2024) is important for maintaining the performance of LLMs after KV cache compression.
-Note that a tradeoff exists: increasing the recent-size allows more infomation to be propagated, while increasing the cache size to be stored.
-To study its impacts, we evaluate the performance produced by three recent-size, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the basleline model.
+Preserving the most recent KV embeddings (Zhang et al., 2024c; Li et al., 2024) is important for maintaining the performance of LLMs after KV cache compression.
+However, a tradeoff exists: increasing the recent-size allows more information to be retained, but also increases the cache size.
+To assess its impact, we evaluate the performance of three recent-size configurations, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the baseline model.
 The results are summarized in Table 6.
-One can observe that a recent-size of 32 yields superior performance than 00 in terms of averaged score on LongBench which demonstrates the importance of keeping the most recent KVs.
-On the other hand, the performance of $32$ and $128$ is negligible, suggesting that maintaining the most recent 32 KVs suffices to preserve optimal performance.
-Figure 5: Performance comparisons under the same KV cache memory consumption.
-Comparison Under the Same Memory Consumption.
-To ensure a fair comparison, we adjust the KV cache size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK.
-For example, the KV cache size of H2O with ThinK is $128$.
-Due to channel pruning of the key cache, the memory consumption of H2O with ThinK at a $128$ KV-size is lower than H2O at a 128 KV-size.
-Therefore, the KV-size of H2O is adjusted from $128$ to $109$.
-Table 6 and Figure 5 present the results of these comparisons on LongBench.
-As the results show, H2O or SnapKV with ThinK deliver superior performance compared to H2O or SnapKV without ThinK while maintaining the same memory usage.
-This demonstrates that KV cache compression methods, combined with our query-based channel pruning, can compress the KV cache more effectively and utilize memory more efficiently.
+As observed, a recent-size of 32 yields superior performance compared to 00, as indicated by the averaged score on LongBench, demonstrating the importance of retaining the most recent KVs.
+On the other hand, the performance difference between recent-sizes of $32$ and $128$ is negligible, suggesting that retaining the most recent 32 KVs is sufficient to maintain optimal performance.
+Figure 3: (a) presents the performance comparison with token eviction methods under identical memory usage for Mistral-7B-Instruct-v0.2, while (b) illustrates the memory usage comparison with the KV cache quantization method KIVI across different batch sizes for LLaMA-2-7B-chat.
+ThinK ($0.4$) indicates we prune the key cache channels with a pruning ratio of $\lambda=0.4$.
+Performance Comparison Under the Same Memory Usage.
+To ensure a fair comparison, we adjust the KV-size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK on Mistral-7B-Instruct-v0.2.
+For example, the KV-size of H2O with ThinK is set to $128$.
+Due to channel pruning applied to the key cache, the memory consumption of H2O with ThinK at a KV-size of $128$ is lower than that of H2O at the same KV-size.
+Consequently, the KV-size of H2O is adjusted from $128$ to $109$ to equalize memory usage.
+Table 7 and Figure 3(a) present the results of these comparisons on the LongBench benchmark.
+The results demonstrate that H2O or SnapKV combined with ThinK consistently outperforms their counterparts without ThinK while maintaining the same memory footprint.
+This highlights the effectiveness of integrating query-driven channel pruning with KV cache compression methods, enabling more efficient memory utilization and improved compression of the KV cache.
+Memory Usage Comparison.
+To evaluate the efficiency of ThinK, we follow the methodology used in KIVI (Liu et al., 2024b).
+We generate synthetic workloads with an input prompt length of 160 and an output length of 338.
+The peak memory usage is reported for the vanilla FP16 baseline, KIVI, and KIVI combined with ThinK ($0.4$) for LLaMA-2-7B-chat.
+As in Figure 3(b), the memory savings from our method become increasingly evident as the batch size grows, in both the KIVI $/2$ 2 and KIVI $/4$ 4 configurations.
+Compared to the baseline model, our approach achieves over a 5$\times$ increase in batch size while maintaining the same memory footprint when integrated with KIVI.
 Pruning Channels of Both Key and Value Cache.
-In our previous study, we evaluate methods with pruning channels of the key cache.
-In this part, we explore the impact of pruning channels of the value cache (Appendix A.1).
-Specifically, for KV cache compression methods, we prune channels of the key cache at one pruning ratio and channels of the value cache at another pruning ratio.
-Table 7 presents the results with LLaMA3-8B and Mistral-7B on LongBench.
-When testing on the base model LLaMA3-8B, H2O or SnapKV with key and value channel pruning can perform on par with H2O or SnapKV without channel pruning.
-In some cases, H2O or SnapKV with key and value channel pruning may even perform better.
-Pruning the channels of the value cache over the base model Mistral-7B experiences a slight performance drop.
-Despite this, note that KV cache with additional value channel pruning can further reduce memory usage.
+In our previous study, we evaluate methods that focused on pruning channels of the key cache.
+In this part, we explore the impact of pruning channels in the value cache as well (Appendix D).
+Specifically, for KV cache compression methods, we apply different pruning ratios to the channels of the key and value caches.
+Table 8 presents the results with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2 on the LongBench benchmark.
+When evaluating the base model LLaMA-3-8B-Instruct, H2O or SnapKV with both key and value channel pruning perform comparably to H2O or SnapKV without pruning.
+In certain instances, the models with key and value channel pruning even outperform their non-pruned counterparts.
+For the base model Mistral-7B-Instruct-v0.2, pruning the value cache channels leads to a slight performance drop.
+Nevertheless, it is important to note that applying value channel pruning to the KV cache further reduces memory usage.
+5 Related Work
+In scenarios involving long contexts, the key-value (KV) cache poses the most significant computational and memory burden within the attention mechanism of large language models.
+Reducing the KV cache is therefore a key priority for optimizing deployment efficiency.
+To address this challenge, system-level optimizations, such as FlashAttention (Dao, 2023) and PagedAttention (Kwon et al., 2023), have been developed to tackle this challenge.
+In parallel, algorithm-level optimizations are also being explored to further enhance efficiency.
+KV Cache Eviction.
+StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with some recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens.
+H2O (Zhang et al., 2024c) selectively retains a small subset of tokens by greedily dropping those with lower contributions to cumulative attention.
+SnapKV (Li et al., 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts.
+FastGen (Ge et al., 2023) adaptively evicts tokens from attention heads that focus on local contexts, discarding non-special tokens that surround key tokens, while standard KV cache is applied to attention heads that attend more broadly.
+PyramidKV (Zhang et al., 2024b) and PyramidInfer (Yang et al., 2024) take a hierarchical approach, adjusting KV cache sizes across different layers by allocating more cache to lower layers and less to higher ones.
+KV Cache Quantization.
+SmoothQuant (Xiao et al., 2023a) enables the quantization of the KV cache to $8$-bit with minimal performance degradation.
+Q-Hitter (Zhang et al., 2024c) leverages accumulated attention scores and ”Quantization Friendliness” metrics to identify tokens that are crucial for preserving the generalization capabilities of LLMs, making them suitable for KV cache quantization.
+Furthermore, recent studies suggest that the key and value caches should be quantized differently (Liu et al., 2024b; Hooper et al., 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
+Structured Pruning of LLMs.
+Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation.
+In contrast, our approach preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache.
+By dynamically identifying unimportant channels using data dependant criterion, our method greatly reduce the key cache size with minimal performance loss.
+6 Conclusion
+Inspired by the observation that certain channels exhibit significantly larger magnitudes than others, and supported by singular value analysis indicating that the key cache is inherently low-rank, we propose ThinK  as a pruning method targeting the key cache channels.
+Our pruning strategy is query-dependent and optimized based on attention scores, ensuring that essential information is preserved for each input query.
+In addition, ThinK can be seamlessly integrated with existing token-level KV cache pruning techniques (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization methods (Liu et al., 2024b), further enhancing inference efficiency.
+Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate the effectiveness our query-dependent channel pruning approach.
+Our method achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%.
+For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.
+Appendix A Observations
+Figure 5 and Figure 4 illustrates the observations which motivates our approach ThinK to prune unimportant key cache channels.
+We conducted singular value decomposition (SVD) (Demmel, 1997) on the attention weights of the specified layers to investigate their principal components.
+Note that
+$\mathbf{U},\mathbf{\Sigma},\mathbf{V}=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}.$
+(a) Key Cache ($14^{th}$ layer)
+(b) Value Cache ($14^{th}$ layer)
+(c) Key Cache ($20^{th}$ layer)
+(d) Value Cache ($20^{th}$ layer)
+Figure 4: Magnitude of key and value cache for LLaMA-2-7B.
+The first head of layer $14$ and layer $20$ of LLaMA-2-7B is selected to visualize the magnitude of the key and value caches.
+We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
+(a) Energy of singular values
+(b) Cumulative energy of singular values
+Figure 5: The energy and cumulative energy of the singular values.
+Appendix B Needle-in-a-Haystack test performance comparison
+Figure 6 visualizes the test performance comparison on Needle-in-a-Haystack on Mistral-7B-Instruct-v0.2.
+(a) KV-size $128$, SnapKV, Acc. $77.8$
+(b) KV-size $128$, SnapKV + ThinK ($0.4$) Acc. $78.6$
+(c) KV-size $1024$, SnapKV, Acc. $90.4$
+(d) KV-size $1024$, SnapKV + ThinK ($0.5$), Acc. $90.8$
+Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$
+Appendix C Implementations
+C.1 Implementation with quantization
+Figure 7 illustrates the implementation of our method when integrated with the KV cache quantization method KIVI (Liu et al., 2024b).
+During the prefill phase, we first prune the unimportant channels of $X_{K}$ before applying quantization.
+In the decoding phase, each newly arrived key cache $t_{K}$ is added to $X_{K_{r}}$.
+Once $X_{K_{r}}$ reaches $G$ tokens, the residual length hyperparameter in KIVI, we prune and quantize the data, then concatenate it with the previously quantized $Q(P(X_{K_{g}}))$.
+Figure 7: Implementations of ThinK when incorporated with KIVI.
+Appendix D Value Cache Pruning
+Similar to the approach used for the Key cache, the pruning of channels in the Value cache can be guided by two primary criteria: magnitude-based pruning and query-driven pruning.
+We find that query-driven pruning is still better than magnitude based pruning.
+$\text{Score}_{v,i}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})[j]=\|\text{softmax}(\frac{\textbf{Q}_{i}[-S_{\text{obs}}:]\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}[:,j]\|_{F}$ (3)
+$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (4)
+where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$.
+We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache.
+Then, only top T channels are retained.
+Table 8 reported the results of pruning both key and value channels.
 Table 8: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench.
-H2O $+$ ThinK($\lambda_{1}$+$\lambda_{2}$) indicates that H2O has $\lambda_{1}$of the Key cache channels pruned and $\lambda_{2}$of the Value cache channels pruned.
+H2O $+$ ThinKV ($\lambda_{1}$+$\lambda_{2}$) indicates that the key cache channels of H2O are pruned with a pruning ratio of $\lambda_{1}$ and the value cache channels are pruned of a pruning ratio of $\lambda_{2}$.
 Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
@@
 RB-P
-LLaMA3 8B KV-size 128
+LLaMA-3-8B-Instruct KV-size 128
 H2O 22.12 13.20 31.61 37.79 32.71 18.45 20.32 22.02 21.10 38.50 87.75 39.14 5.83 69.50 55.06 50.97 35.38
-+ThinK(30+30) 23.71 13.65 33.08 41.86 29.88 18.04 19.60 21.65 20.26 38.00 86.08 38.61 5.16 69.50 57.59 55.19 35.74
++ThinKV (0.3+0.3) 23.71 13.65 33.08 41.86 29.88 18.04 19.60 21.65 20.26 38.00 86.08 38.61 5.16 69.50 57.59 55.19 35.74
 […]
-+ThinK(30+30) 21.86 13.79 33.26 40.93 29.39 19.22 18.81 21.30 19.26 41.50 87.00 37.95 5.78 69.50 57.84 55.62 35.81
++ThinK(0.3+0.3) 21.86 13.79 33.26 40.93 29.39 19.22 18.81 21.30 19.26 41.50 87.00 37.95 5.78 69.50 57.84 55.62 35.81
 KV-size 512
 H2O 23.52 17.93 34.68 42.11 33.52 19.92 22.11 22.56 23.82 41.00 90.46 40.20 5.87 69.50 56.71 51.69 37.23
-+ThinK(30+30) 22.83 17.57 34.18 42.67 33.52 19.95 21.17 22.23 22.82 38.50 90.11 39.08 5.21 69.0 59.99 56.83 37.23
++ThinKV (0.3+0.3) 22.83 17.57 34.18 42.67 33.52 19.95 21.17 22.23 22.82 38.50 90.11 39.08 5.21 69.0 59.99 56.83 37.23
 […]
-+ThinK(30+30) 24.57 24.59 38.09 44.61 34.37 20.37 21.23 21.95 23.30 66.00 90.69 39.38 5.60 69.00 61.75 58.46 40.25
++ThinKV (0.3+0.3) 24.57 24.59 38.09 44.61 34.37 20.37 21.23 21.95 23.30 66.00 90.69 39.38 5.60 69.00 61.75 58.46 40.25
 KV-size 2048
 H2O 25.56 26.85 39.54 44.30 32.92 21.09 24.68 23.01 26.16 53.00 90.65 41.84 4.91 69.25 58.43 51.31 39.59
-+ThinK(30+30) 25.03 26.77 39.68 42.12 33.08 19.59 23.00 22.89 25.27 51.00 91.11 40.58 5.23 69.00 61.12 57.95 39.59
-+ThinK(40+40) 24.87 24.31 37.77 43.13 34.42 19.60 21.67 22.70 24.52 49.00 90.81 39.28 6.00 69.00 61.81 58.08 39.19
++ThinKV (0.3+0.3) 25.03 26.77 39.68 42.12 33.08 19.59 23.00 22.89 25.27 51.00 91.11 40.58 5.23 69.00 61.12 57.95 39.59
++ThinKV (0.4+0.4) 24.87 24.31 37.77 43.13 34.42 19.60 21.67 22.70 24.52 49.00 90.81 39.28 6.00 69.00 61.81 58.08 39.19
 […]
-+ThinK(30+30) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
-+ThinK(40+40) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
-Mistral 7B KV-size 128
++ThinKV (0.3+0.3) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
++ThinKV (0.4+0.4) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
+Mistral-7B-Instruct-v0.2 KV-size 128
 H2O 21.21 21.81 38.87 30.42 20.36 12.30 20.58 22.61 22.10 39.00 82.37 40.44 2.90 79.56 51.22 48.38 34.63
-+ThinK(30+30) 20.71 21.49 38.01 30.66 22.28 13.87 20.13 22.45 21.07 38.50 82.20 38.69 2.94 78.56 51.55 48.28 34.46
++ThinKV (0.3+0.3) 20.71 21.49 38.01 30.66 22.28 13.87 20.13 22.45 21.07 38.50 82.20 38.69 2.94 78.56 51.55 48.28 34.46
 […]
-+ThinK(30+30) 19.92 20.61 42.68 37.63 23.19 15.09 18.97 21.93 20.55 45.00 84.06 39.33 2.99 66.00 51.51 47.51 34.81
++ThinKV (0.3+0.3) 19.92 20.61 42.68 37.63 23.19 15.09 18.97 21.93 20.55 45.00 84.06 39.33 2.99 66.00 51.51 47.51 34.81
 KV-size 512
 H2O 21.83 26.00 44.69 32.46 23.05 14.69 23.53 23.06 24.59 42.00 85.22 41.49 3.40 86.20 54.78 51.09 37.38
-+ThinK(30+30) 22.36 24.26 44.77 30.47 22.94 14.96 22.63 22.90 23.73 41.50 85.30 40.21 3.08 80.07 54.48 50.96 36.54
-+ThinK(30+10) 22.14 25.15 45.29 31.78 23.21 14.62 23.36 22.70 24.51 41.50 85.61 41.58 2.75 84.03 54.50 51.09 37.11
++ThinKV (0.3+0.3) 22.36 24.26 44.77 30.47 22.94 14.96 22.63 22.90 23.73 41.50 85.30 40.21 3.08 80.07 54.48 50.96 36.54
++ThinKV (0.3+0.1) 22.14 25.15 45.29 31.78 23.21 14.62 23.36 22.70 24.51 41.50 85.61 41.58 2.75 84.03 54.50 51.09 37.11
 […]
-+ThinK(30+30) 24.10 27.04 47.76 38.66 25.45 17.51 22.64 22.81 23.91 66.00 86.62 39.91 3.36 82.24 55.96 52.81 39.80
-+ThinK(30+10) 23.90 28.14 48.35 39.03 24.83 16.68 23.51 23.12 24.34 67.50 86.09 41.69 2.65 84.34 57.29 53.22 40.29
++ThinKV (0.3+0.3) 24.10 27.04 47.76 38.66 25.45 17.51 22.64 22.81 23.91 66.00 86.62 39.91 3.36 82.24 55.96 52.81 39.80
++ThinKV (0.3+0.1) 23.90 28.14 48.35 39.03 24.83 16.68 23.51 23.12 24.34 67.50 86.09 41.69 2.65 84.34 57.29 53.22 40.29
 KV-size 1024
 H2O 23.67 28.55 46.4 36.99 24.82 15.02 25.21 23.04 25.77 46.00 85.93 41.98 3.24 86.57 56.40 52.75 38.90
-+ThinK(30+30) 23.65 26.54 47.00 35.52 24.79 17.15 23.64 23.12 25.20 44.00 86.38 41.67 3.46 80.14 56.53 52.86 38.23
-+ThinK(30+10) 24.13 28.57 46.31 35.59 24.92 15.34 24.58 23.33 25.93 45.50 85.91 42.97 2.57 83.64 55.39 52.73 38.59
++ThinKV (0.3+0.3) 23.65 26.54 47.00 35.52 24.79 17.15 23.64 23.12 25.20 44.00 86.38 41.67 3.46 80.14 56.53 52.86 38.23
++ThinKV (0.3+0.1) 24.13 28.57 46.31 35.59 24.92 15.34 24.58 23.33 25.93 45.50 85.91 42.97 2.57 83.64 55.39 52.73 38.59
 […]
-+ThinK(30+30) 25.29 29.25 49.17 41.25 25.75 19.37 24.64 23.02 25.27 69.00 86.70 40.92 3.29 82.06 57.15 54.15 41.02
-+ThinK(30+10) 25.84 29.30 49.56 41.44 25.29 19.02 25.21 23.73 25.72 69.00 86.69 42.55 2.44 85.76 57.55 54.10 41.45
++ThinKV (0.3+0.3) 25.29 29.25 49.17 41.25 25.75 19.37 24.64 23.02 25.27 69.00 86.70 40.92 3.29 82.06 57.15 54.15 41.02
++ThinKV (0.3+0.1) 25.84 29.30 49.56 41.44 25.29 19.02 25.21 23.73 25.72 69.00 86.69 42.55 2.44 85.76 57.55 54.10 41.45
 KV-size 2048
 H2O 25.76 31.10 49.06 40.38 26.43 16.78 27.17 23.64 26.69 55.0 86.35 42.48 2.72 86.64 56.98 53.91 40.69
-+ThinK(30+30) 25.60 28.74 47.54 38.67 26.25 17.35 24.54 23.27 26.15 51.00 87.01 43.02 2.94 81.46 56.41 54.26 39.64
-+ThinK(30+10) 25.64 30.65 48.95 40.42 26.43 16.65 26.76 23.51 26.59 52.50 86.53 43.45 2.66 83.96 56.55 53.83 40.32
++ThinKV (0.3+0.3) 25.60 28.74 47.54 38.67 26.25 17.35 24.54 23.27 26.15 51.00 87.01 43.02 2.94 81.46 56.41 54.26 39.64
++ThinKV (0.3+0.1) 25.64 30.65 48.95 40.42 26.43 16.65 26.76 23.51 26.59 52.50 86.53 43.45 2.66 83.96 56.55 53.83 40.32
 […]
-+ThinK(30+30) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
-+ThinK(30+10) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94
-(a) KV-size 128, SnapKV, Acc.77.8
-(b) KV-size 128, SnapKV+ThinK($40$) Acc.78.6
-(c) KV-size 1024, SnapKV, Acc.90.4
-(d) KV-size 1024, SnapKV+ThinK($50$), Acc.90.8
-Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2.
-5 Related Work
-In scenarios involving long contexts, the most significant computational burden from the attention mechanism is the key-value (KV) cache.
-Reducing the KV cache is a high priority for optimizing deployment efficiency.
-System-level optimizations, such as FlashAttention (Dao,, 2023) and PagedAttention (Kwon et al.,, 2023), have been developed to address this issue.
-Additionally, algorithm-level optimizations are being explored to further enhance efficiency.
-KV Cache Eviction.
-StreamingLLM (Xiao et al., 2023b, ) maintains a few initial tokens and some recent tokens based on the observation of attention sink, which may result in the loss of important information carried by the dropped tokens.
-H2O (Zhang et al., 2024c, ) retains only a small portion of the tokens by greedily dropping tokens based on their contributions to the cumulative attention.
-SnapKV (Li et al.,, 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts.
-FastGen (Ge et al.,, 2023) adaptively evicts tokens from attention heads that emphasize local contexts.
-This approach focuses on discarding non-special tokens centered on special tokens, while the standard KV cache is used only for attention heads that broadly attend to all tokens.
-PyramidKV (Zhang et al., 2024b, ) and PyramidInfer (Yang et al.,, 2024) considers adjusting the KV cache size across different layers by allocating more cache in the lower layers and less in the higher ones.
-KV Cache Quantization.
-SmoothQuant (Xiao et al., 2023a, ) can quantize the KV cache to 8-bit with minimal performance degradation.
-Q-Hitter (Zhang et al., 2024c, ) uses accumulated attention scores and ‘Quantization Friendliness’ metrics to identify tokens that are essential for maintaining the generalization capabilities of LLMs and are suitable for KV cache quantization.
-Some studies have found that the key cache and value cache should be quantized differently (Liu et al., 2024b, ; Hooper et al.,, 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
-Structured Pruning of LLMs.
-Structured pruning  (Ma et al.,, 2023; Ding et al.,, 2023) of LLMs traditionally focuses on removing unimportant layers, heads, and hidden dimensions, which often results in significant performance degradation.
-In contrast, our methodology preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache.
-By dynamically identifying unimportant channels based on data dependant criterion, our approach greatly reduce the key cache-size with negligible performance loss.
-6 End Note and Future Direction
-Motivated by the observation that certain channels have significantly larger magnitudes compared to others, and the singular value analysis indicates that the key cache is inherently low-rank (Singhania et al.,, 2024), we propose ThinK to perform pruning over the key cache channel.
-The proposed pruning strategy is query-dependant and optimized based on the attention scores, ensuring that essential information is retained for each input query.
-In addition, ThinK can be seamlessly integrated with other popular token-level KV cache quantization techniques (Li et al.,, 2024; Zhang et al., 2024c, ), further enhancing inference efficiency.
-Extensive experiments on LongBench (Bai et al.,, 2023) and Needle-in-a-Haystack tests with two foundation models demonstrate the effectiveness and robustness of our query-dependent channel pruning method.
-Our approach achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%.
-Our analysis indicates that ThinK can maintain superior performance over baselines with a smaller KV cache size under equivalent memory consumption conditions.
-Future work will focus on strategies to increase the pruning ratio without compromising model performance, and exploring pruning techniques for value caches.
-We also plan to evaluate channel importance from a more sophisticated and compositional perspective, incorporating both token-level and channel-level information to calculate importance.
-Appendix A Additional Experimental Details
-A.1 Value Cache Pruning
-Similar to the approach used for the Key cache, the pruning of channels in the Value cache can be guided by two primary criteria: magnitude-based pruning and query-driven pruning.
-We find that query-driven pruning is still better than magnitude based pruning.
-$\text{Score}_{v,i}(\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i})[j]=\|\text{softmax}(\frac{\textbf{Q}_{i}[-S_{\text{obs}}:]\textbf{K}_{i}^{T}}{\sqrt{D}})\textbf{V}_{i}[:,j]\|_{F}$ (10)
-$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (11)
-where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$.
-We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache.
-Then, only top T channels are retained.
++ThinKV (0.3+0.3) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
++ThinKV (0.3+0.1) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94

### v3

@@
 Our approach not only maintains or enhances model accuracy but also achieves a reduction in KV cache memory costs by over 20% compared with vanilla KV cache eviction and quantization methods.
-For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.
-Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK, establishing a new baseline algorithm for efficient LLM deployment without compromising performance.
+For instance, ThinK integrated with KIVI can achieve 2.8×2.8$\times$2.8 × peak memory reduction while maintaining nearly the same quality, enabling a batch size increase from 4$\times$ (with KIVI alone) to 5$\times$ when using a single GPU.
+Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK.
+Our code has been made available at https://github.com/SalesforceAIResearch/ThinK.
 1 Introduction
@@
 Consequently, efficient LLMs have gained traction in recent years (Hu et al., 2021; Wan et al., 2023).
-To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) are employed to reduce model size.
+To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023; Lu et al., 2024b) are employed to reduce model size.
 Additionally, the key-value (KV) cache, stored in GPU memory alongside model parameters, scales linearly with both sequence length and batch size, creating a substantial memory burden when handling long sequences.
@@
 Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows: $2\times B\times S\times L\times N\times D$, where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension.
-The KV cach size grows linearly as the batch size $B$ and sequence length $S$.
+The KV cache size grows linearly as the batch size $B$ and sequence length $S$.
 For a model with multihead attention (MHA) (Vaswani et al., 2017), such as LLaMA2-7B (Touvron et al., 2023b), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters.
@@
 Magnitude based Pruning: Based on the observations in Figure 4 which depicts the significant variation in the magnitudes across different channels, one straightforward criterion is to use the norm of the magnitude to measure the importance of different channels in key cache.
-$M_{n,d}=\big{\|}K[n,:,d]\big{\|}_{p}.$ (1)
+$M_{n,d}=\big{\|}\textbf{K}[n,:,d]\big{\|}_{p}.$ (1)
 Given pruning ratio $\lambda$, We only keep $T=\lfloor(1-\lambda)D\rfloor$ most important channels among the $D$ channels of each head: $I=\textbf{Top}_{T}(M,T)$ where $\|\cdot\|_{p}$ is the $l_{p}$ norm of each channel.
@@
 The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension.
-These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.
+These results also indicate the need for a better pruning metrics to achieve higher pruning ratios effectively.
 3.2 Query-Driven Pruning
@@
 Observation Window.
-Based on the observations in SnapKV (Li et al., 2024) that the last window of input sequence recognizes highly similar attention pattern with generation.
-To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
+Following SnapKV (Li et al., 2024), to reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score as the last window of input sequence recognizes highly similar attention pattern with generation: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
 3.3 Implementation of ThinK
 Figure 2: Implementation during decoding.
-Following SnapKV, we focus on the long context input scenario.
 We opt not to prune the most recent tokens and newly generated keys.
-Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size.
-Additionally, we maintain a binary mask to indicate which channels have been pruned.
-it is important to note that the memory overhead associated with this mask is negligible.
+Consequently, our key-value (KV) cache has two categories: one subset consists of pruned keys with a reduced channel size, while the other at their original size.
+Additionally, we maintain a binary mask whose memory overhead is negligible to indicate which channels have been pruned.
 Figure 2 illustrates one implementation of ThinK during the decoding stage.
-This implementation begins by pruning the query using the mask.
+We first prune the query using the stored mask, ensuring that the query dimensionality and the pruned key cache remain consistent.
 The pruned query is then multiplied by the pruned key, while the unpruned query is applied to the unpruned key.
 Subsequently, the two outputs are concatenated.
+Our method removes unimportant Key cache channels to get real memory reduction.
 Table 2: Performance comparison of key cache pruning on LLaMA-3-(8B/70B)-Instruct on LongBench. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$.
@@
 In this section, we conduct comprehensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction.
-First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1.
-Next, we report the performance ThinK with two KV eviction methods and one KV quantization methods in long-context scenarios in Section 4.2 and  4.3.
-Finally, we present the results of ablation experiments in Section 4.4.
+In the experiments, we prune Key cache channels by default, the value cache results are in Table 8 in Appendix D.
 4.1 Settings
@@
 For instance, when comparing SnapKV and SnapKV integrated with ThinK, we used a maximum pooling kernel size of 7 and an observation window size of 32, maintaining the same KV-size for both configurations.
+We compress the Key cache starting from the prefilling stage.
 Table 4: Performance evaluation of combining ThinK with KIVI Liu et al. (2024b) on LongBench. ThinK ($0.4$) indicates we prune the key cache channels with a pruning ratio of $\lambda=0.4$.
@@
 Tables 2 and  3 present the results of KV compression methods and their integration with our proposed channel pruning technique for the key cache (ThinK) across three different base LLMs, evaluated at various KV-sizes on the LongBench benchmark.
+The pruning ratio of $\lambda=0.4$ indicate that 40% of key cache channels are removed ,resulting in a 20% reduction in the total KV cache memory footprint.
 The following observations can be drawn: (1) Our method successfully prunes the channels of the key cache after the KV cache has been compressed using H2O and SnapKV.
@@
 Moreover, query-driven channel pruning proves more effective than $l_{1}$ and $l_{2}$ norm-based methods for channel pruning in LLMs.
+Experiment results of applying ThinK directly to vanilla models are in Table 9 which further validate the effectiveness of our method.
 We further validate the efficacy of our method by applying it to the KV cache quantization technique KIVI (Liu et al., 2024b), as shown in Table 4.
@@
 In addition, the performance on larger KV-sizes suggests that ThinK is robust for long-context tasks.
-Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depth percentages.
+Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depths.
 The KV-sizes are set to $128$ and $1024$, with pruning ratios of $\lambda=0.4$ and $\lambda=0.5$, respectively.
@@
 As in Figure 3(b), the memory savings from our method become increasingly evident as the batch size grows, in both the KIVI $/2$ 2 and KIVI $/4$ 4 configurations.
-Compared to the baseline model, our approach achieves over a 5$\times$ increase in batch size while maintaining the same memory footprint when integrated with KIVI.
+Compared to the baseline model, our approach achieves over a 5$\times$ (from 4$\times$ with KIVI alone) increase in batch size while maintaining the same memory footprint when integrated with KIVI.
+Model weights and KV cache are the primary memory components accessed during generation.
+By effectively reducing the memory footprint of the KV cache, our method alleviates the memory bottleneck, enabling faster generation speeds as shown in Table 10.
 Pruning Channels of Both Key and Value Cache.
-In our previous study, we evaluate methods that focused on pruning channels of the key cache.
-In this part, we explore the impact of pruning channels in the value cache as well (Appendix D).
+In this part, we explore the impact of pruning channels in the value cache (Appendix D).
 Specifically, for KV cache compression methods, we apply different pruning ratios to the channels of the key and value caches.
@@
 For the base model Mistral-7B-Instruct-v0.2, pruning the value cache channels leads to a slight performance drop.
-Nevertheless, it is important to note that applying value channel pruning to the KV cache further reduces memory usage.
+This aligns with the observations in Figure 4, where the Key cache shows highly unbalanced magnitudes along the channel dimension, while the Value cache exhibits more uniform magnitudes.
+This suggests that the Value cache has less channel sparsity, making it harder to identify redundant channels for pruning.
+Nevertheless, pruning Value cache channels still contributes to further memory reduction in the KV cache.
 5 Related Work
@@
 KV Cache Eviction.
-StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with some recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens.
+StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens.
 H2O (Zhang et al., 2024c) selectively retains a small subset of tokens by greedily dropping those with lower contributions to cumulative attention.
@@
 Structured Pruning of LLMs.
-Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation.
+Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023; Lu et al., 2024a) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation.
 In contrast, our approach preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache.
 By dynamically identifying unimportant channels using data dependant criterion, our method greatly reduce the key cache size with minimal performance loss.
-6 Conclusion
-Inspired by the observation that certain channels exhibit significantly larger magnitudes than others, and supported by singular value analysis indicating that the key cache is inherently low-rank, we propose ThinK  as a pruning method targeting the key cache channels.
-Our pruning strategy is query-dependent and optimized based on attention scores, ensuring that essential information is preserved for each input query.
-In addition, ThinK can be seamlessly integrated with existing token-level KV cache pruning techniques (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization methods (Liu et al., 2024b), further enhancing inference efficiency.
-Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate the effectiveness our query-dependent channel pruning approach.
-Our method achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%.
-For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.
+6 Conclusion and Limitations
+Motivated by the observation that certain channels exhibit significantly larger magnitudes and the low-rank nature of the key cache (as shown by singular value analysis), we propose ThinK as a query-dependent pruning method targeting key cache channels.
+Optimized based on attention scores, our approach preserves essential information for each input query.
+ThinK seamlessly integrates with existing token-level KV cache pruning (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization (Liu et al., 2024b), further improving inference efficiency.
+Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate its effectiveness, achieving comparable or superior performance to baselines while reducing key cache size by 40%.
+However, our method slightly increases TTFT (Time-to-First-Token) due to channel importance computation (see Appendix F).
+Additionally, while effective for key cache pruning, value cache pruning remains an open direction for future work.
 Appendix A Observations
@@
 Then, only top T channels are retained.
-Table 8 reported the results of pruning both key and value channels.
+Table 8 reported the results of pruning both key and value channels, showing that pruning the Value cache channels is harder than pruning the Key cache channels.
+However, pruning 30% of both the Key cache and Value cache on LLaMA-3-8B-Instruct still results in acceptable performance.
+This demonstrates that the Value cache also has potential for pruning in the channel dimension.
+As depicted in Figure 4, the magnitude of the Key cache along the channel dimension is highly unbalanced, whereas the magnitude of the Value cache along the channel dimension is more uniform.
+This indicates that the channel sparsity in the Value cache is not as significant as in the Key cache, making it more challenging to identify redundant channels for pruning.
+We will investigate the value cache pruning strategy as part of our future work.
 Table 8: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench.
@@
 +ThinKV (0.3+0.1) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94
+Appendix E Pruning Key Cache on Vanilla Models
+To furthur demonstrate the effectiveness of our proposed ThinK, we conducted additional experiments applying ThinK directly to vanilla models, specifically LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2.
+The results, as shown in Table 9, demonstrate that our method maintains superior performance even after pruning 40% of the key cache channels.
+Furthermore, when the pruning ratio is increased to 50%, the performance degradation remains within an acceptable range.
+These findings further validate the effectiveness of ThinK, not only for pruned models but also when applied directly to vanilla models.
+Table 9: Performance comparison of pruning key cache on vanilla models on LongBench.
+Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
+NrtvQA
+Qasper
+MF-en
+HotpotQA
+2WikiMQA
+Musique
+GovReport
+QMSum
+MultiNews
+TREC
+TriviaQA
+SAMSum
+PCount
+PRe
+Lcc
+RB-P
+LLaMA-3-8B-Instruct, KV-size Full
+Vanilla 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
++ThinK(0.4) 25.32 32.26 39.81 44.19 34.77 21.10 28.63 23.13 26.38 73.50 90.58 41.69 5.21 69.50 61.94 58.37 42.27
++ThinK(0.5) 25.35 32.80 41.64 43.99 31.81 21.58 28.35 23.31 26.80 73.50 90.37 40.81 5.67 69.17 61.90 59.00 42.25
++ThinK(0.6) 24.39 31.16 41.80 42.52 31.63 20.70 25.84 23.18 25.46 73.50 90.43 38.97 5.77 68.46 59.63 59.38 41.43
+Mistral-7B-Instruct-v0.2, KV-size Full
+Vanilla 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
++ThinK(0.4) 27.11 33.46 48.73 41.79 28.14 18.87 32.45 24.55 27.09 71.00 86.26 43.02 3.95 86.31 56.99 54.36 42.76
++ThinK(0.5) 26.63 33.71 49.38 42.38 26.78 18.76 32.57 24.63 26.92 71.00 86.39 42.82 3.13 84.65 56.75 54.04 42.53
++ThinK(0.6) 27.03 33.23 49.49 42.65 26.43 18.14 31.74 24.75 26.57 71.00 86.28 41.40 3.50 83.11 55.93 53.37 42.16
+Appendix F Comparisons of Generation Speed and Throughput
+In this section, we provide comparisons of TTFT (Time To First Token), TPOT (Time Per Output Token), memory usage, and throughput in Table 10 and Table 11.
+We follow the methodology used in KIVI Liu et al.
+(2024b).
+We generate synthetic workloads with an input prompt length of $160$ and an output length of $338$.
+We set a batch size 300 for both KIVI and our method.
+As our method performs online pruning, there is a slight delay introduced during the prefilling stage due to the computation of channel importance.
+However, the reduction in memory usage leads to notable improvements in TPOT during inference.
+Specifically, after applying ThinK, TPOT improves from 0.27 ms/token to 0.25 ms/token with 40% pruning, and further improves to 0.24 ms/token with a 50% pruning ratio.
+Regarding throughput, with the same memory usage, ThinK (40% pruning) improves the throughput of KIVI from 5168 tokens/s to 5518 tokens/s.
+At a pruning ratio of 50%, throughput increases further to 5676 tokens/s.
+Model weights and KV cache are the primary memory components accessed during generation.
+By effectively reducing the memory footprint of the KV cache, our method alleviates the memory bottleneck, enabling larger batch sizes and faster generation speeds.
+These results highlight the potential of ThinK to enhance inference performance by balancing memory efficiency with computational overhead.
+We are continuing to optimize the implementation to further improve performance.
+Memory bandwidth is a major performance bottleneck in the decoding phase of large language models (LLMs).
+Table 10: Comparisons of TTFT (Time To First Token), TPOT (Time Per Output Token) and memory usage on LLaMA-2-7B.
+Method KIVI(4/4) KIVI(4/4) + ThinK(0.4) KIVI (4/4) + ThinK(0.5)
+Memory (GB) 61.7 53.3 51.2
+TTFT (ms) 7.0 10.4 10.2
+TPOT (ms/token) 0.27 0.25 0.24
+Table 11: Comparisons of throughput on LLaMA-2-7B.
+Method Vanilla KIVI(4/4) KIVI(4/4) + ThinK(0.4) KIVI (4/4) + ThinK(0.5)
+Throughput (tokens/s) 2557 5168 5518 5676
+Appendix G Comparisons with SVD based Methods
+In this section, we compare our ThinK with SVD based methods (Saxena et al., 2024; Yuan et al., 2023; Chang et al., 2024).
+SVD based KV cache compression method decompose the KV cache weights offline with some calibration data, relying on storing latent representations and recovering the cache during inference.
+Our approach is an online pruning strategy that is plug-and-play, requiring no changes to the model’s weights or architecture.
+This makes ThinK lightweight and easy to integrate into existing systems.
+Besides, our method dynamically prunes the Key cache channels directly without requiring any reconstruction.
+We provide comparisons of our method with SVD-based approaches in Table 12.
+Following Palu, we use Mistral-7B-Instruct-v0.2 as the baseline model.
+Our results show that when the KV cache is compressed by 80%, our method preserves performance, whereas Palu experiences an accuracy drop of 1.28%.
+Similarly, for ASVD, we evaluate under the same compression rate.
+Our method demonstrates significantly better performance, with only a 0.12% performance drop compared to the 5.21% drop observed for ASVD.
+The results demonstrate the effectiveness of ThinK, with significantly less performance degradation compared to SVD-based methods under the same compression rates.
+Our pruning method has the potential to be integrated with SVD-based approaches.
+Exploring such a combination could yield further advancements in KV cache compression, which we aim to investigate in future work.
+Table 12: Performance comparison of SVD based methods on LongBench.
+Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
+NrtvQA
+Qasper
+MF-en
+HotpotQA
+2WikiMQA
+Musique
+GovReport
+QMSum
+MultiNews
+TREC
+TriviaQA
+SAMSum
+PCount
+PRe
+Lcc
+RB-P
+Mistral-7B-Instruct-v0.2
+Vanilla 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
++Palu 27.47 34.11 48.47 44.09 26.33 20.18 31.15 24.30 27.12 70.00 85.80 41.74 2.74 73.18 51.70 54.52 41.43
++ThinK(0.4) 27.11 33.46 48.73 41.79 28.14 18.87 32.45 24.55 27.09 71.00 86.26 43.02 3.95 86.31 56.99 54.36 42.76
+LLaMA-2-7B-Chat
+Vanilla 18.39 20.11 35.67 31.25 25.50 10.14 25.68 20.93 26.27 64.00 83.38 40.99 5.50 10.00 60.81 55.27 33.37
++ASVD 16.46 13.19 28.98 21.94 22.86 8.74 17.73 20.27 22.35 57.00 73.88 38.51 1.50 4.77 53.32 49.13 28.16
++ThinK(0.4) 18.39 19.98 35.05 30.85 25.65 10.25 25.98 20.82 26.04 64.00 83.63 41.55 6.00 8.50 60.18 55.18 33.25